### v1

# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

Abstract
  1. From common-sense reasoning to domain-specific tasks, parameter-efficient fine tuning (PEFT) methods for large language models (LLMs) have showcased significant performance improvements on downstream tasks. However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions, particularly due to sparse training data. This overconfidence reflects poor epistemic uncertainty calibration, which arises from limitations in the model’s ability to generalize with limited data. Existing PEFT uncertainty quantification methods for LLMs focus on the post fine-tuning stage and thus have limited capability in calibrating epistemic uncertainty. To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates functional-level epistemic uncertainty during the fine-tuning stage via a mixture-of-expert framework. We show that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks. Furthermore, UQ4CT maintains superior ECE performance with high accuracy under distribution shift, showcasing improved generalizability.
11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint. Under Review. Correspondence to: Yian Ma <yianma@ucsd.edu>.
1 Introduction
Large Language Models (LLMs) have revolutionized various domains as general task solvers (Chang et al., 2024). To adapt LLMs for specific downstream tasks or create instruction-following models, fine-tuning have become increasingly important (Houlsby et al., 2019; Hu et al., 2021a; Liu et al., 2022; Ding et al., 2022, 2023). This involves additional training on pre-trained LLMs using a smaller dataset (Zhong et al., 2021; Ren et al., 2022). Through fine-tuning, the model parameters are updated to better adapt to the domain-specific knowledge (Peng et al., 2023). To reduce the computational cost for fine-tuning, Hu et al. (2021a) proposed Low-Rank Adaptation (LoRA), which effectively reduces the parameters required for fine-tuning by introducing low-rank trainable matrices at each layer of the transformer architecture instead of fine-tuning the full model parameters. Li et al. (2024); Wu et al. (2024a) proposed LoRA Mixture-of-Experts (MoE) models which grants better performance while maintaining parameter efficiency.
However, previous studies have shown that fine-tuned LLMs are often overconfident with their predictions (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023). This resembles poorly calibrated uncertainty (Zhou et al., 2022) due to the sparsity of fine-tuning data. Overconfidence is a crucial problem in safety-related decision making or in fields where data is very limited, such as experimental design, climate science and medical diagnosis (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022). Thus, methods that enhance uncertainty quantification of fine-tuned LLMs is urgently needed to assure trustworthy predictions for better application.
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure. Monte-Carlo dropout (Gal and Ghahramani, 2016a) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023a) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates. Although these methods have demonstrated improved uncertainty estimations, they all utilize a fixed set of LoRA parameters fine-tuned over the entire downstream task dataset. The point estimates of parameters have very limited capabilities in capturing epistemic uncertainty, while direct calibration of epistemic uncertainty over the entire LoRA parameter space is an ideal but not practical approach.
Therefore, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the functional-level epistemic uncertainty via the LoRA MoE architecture. We propose a functional perspective on LoRA parameters where we treat the LoRA experts as basis functions and consider the more complex, prompt dependent functions as mixtures of those basis functions. On top of learning the parameters in the LoRA experts, UQ4CT also trains a prompt-dependent LoRA mixture to form a calibrated distribution over the functional space. The LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input. The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space. We calibrate functional level epistemic uncertainty to align with predictive correctness during training time. This significantly improves uncertainty estimations of the model on its predictions without compromising the accuracy. To summarize, our contributions include:
  • A novel epistemic uncertainty quantification approach with Mixture-of-Experts architecture during fine-tuning stage to model functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
  • A novel training calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
  • More than $25\%$ Expected Calibration Error reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task, superior ECE performance under distribution shift scenarios on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
2 Preliminaries
2.1 Low-rank Adaptation (LoRA)
LLMs have numerous large weight matrices to perform matrix multiplication, denoted as $\mathbf{W}_{0}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}}$ that maps inputs $\mathbf{x}$ to outputs $\mathbf{h}$. Hu et al. (2021a) proposes LoRA, which fixes $\mathbf{W}_{0}$ and introduces a low-rank perturbation $\Delta\mathbf{W}$ to the weight matrix:
$\mathbf{h}$ $=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (1)
Here, $\Delta\mathbf{W}$ is calculated as the product of two matrices, $\mathbf{B}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{lr}}}$ and $\mathbf{A}\in\mathbb{R}^{n_{\text{lr}}\times n_{\text{in}}}$ where $n_{\text{lr}}$ is significantly smaller than $n_{\text{in}}$ or $n_{\text{out}}$. For example, we use $n_{\text{lr}}=32$ while $n_{\text{in}}=n_{\text{out}}=4096$ for the Llama2-7b model (Touvron et al., 2023a). Therefore, the total number of LoRA parameters for this $\Delta\mathbf{W}$ is $n_{\text{lr}}(n_{\text{in}}+n_{\text{out}})$, which is far smaller than the parameter count of the full matrix, $n_{\text{in}}n_{\text{out}}$. One of the key motivations of incorporating LoRA to fine-tune LLMs is the vast amount of memory cost reduction compared with fine-tuning on the full model. For an LLM with $7$ billion parameters, maintaining the average gradient and average squared gradients for optimization multiplies the memory required by a factor of $3$ compared to simply loading model weights. LoRA greatly mitigates this memory cost as the tripled memory consumption only applies to LoRA adapters.
2.2 Mixture of Experts (MoE)
LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is a efficient approach to scale the number of parameters while maintaining the same computational bounds. LoRA MoE utilizes the top-k router to assign each token to the LoRA experts (Lepikhin et al., 2020). The router is a linear layer that maps the input hidden state $\mathbf{h}$ to a probability distribution of candidate experts.
Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router chooses $k$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-k}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (2)
Finally, we obtain the final MixLoRA prediction with:
$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{i}\mathbf{A}^{\ell}_{i}\cdot\mathbf{h}^{\ell}$ (3)
where $\mathbf{W}$ is the pretrained weights of the feed-forward network (FFN) layer and $\mathbf{B}^{\ell}_{i}\mathbf{A}^{\ell}_{i}$ is the i-th LoRA expert.
2.3 Aleatoric and Epistemic Uncertainties
In machine learning models, uncertainty can be categorized into aleatoric (data-wise) and epistemic (model-wise) uncertainty (Hora, 1996; Hüllermeier and Waegeman, 2021). For LLMs, aleatoric uncertainty arises from the inherently ambiguous and context dependent nature of natural languages where a single phrase or sentence can have multiple valid interpretations in different contexts. Epistemic uncertainty is introduced by the model’s lack of knowledge due to limited learning capabilities, suboptimal modeling or sparse training data.
Epistemic uncertainty is highly related to several well-known limitations of generative models. For example, it has been observed that when an LLM is pretrained on a diverse range of text data, it is generally well-calibrated, i.e. the predicted probability of the next token generally aligns with what is observed in real text. However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a). A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target downstream task (Lin et al., 2023; Luo et al., 2023).
Motivated by these observations, we explore functional-level epistemic uncertainty in generative models and aim to develop metrics that assess model performance on specific problem instances to fine-tune the parameter mixture.
3 Methodology
The high level goal of UQ4CT is to balance the exploration and exploitation of different LoRA experts during fine-tuning. In particular, we incorporate the functional-level epistemic uncertainty (FEU) to calibrate the prompt-dependent parameter mixture with LoRA MoE.
Assume that our answer $a$ is generated via a mixture of mechanisms or models $M$, conditioning on the input prompt $x$. Assume that $e(a)$ is an embedding of $a$ so that least squares distance is a natural distance on the space of $e(a)$. For an expressive enough model class $\mathcal{M}$, and a calibrated distribution $P(M|x)$ over the model class, we can measure the deviation of the generated answer to the “ideal” one $a_{*}=f_{*}(x)$ as:
$\mathbb{E}_{M\sim P(M|x)}\left[\mathbb{E}_{a\sim P(a|M,x)}\left[\|e(a)-e(a_{*})\|^{2}\right]\right]$
$=\mathbb{E}_{P(M|x)}\underbrace{\left[\mathbb{E}_{P(a|M,x)}\left[\|e(a)-\mathbb{E}_{P(a|M,x)}[e(a)]\|^{2}\right]\right]}_{\text{Aleatoric Uncertainty}}+\underbrace{\mathbb{E}_{P(M|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-e(a_{*})\|^{2}\right]}_{\text{Epistemic Uncertainty}}.$ (4)
Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$. We hence name the task “functional-level uncertainty quatification”.
3.1 Functional-level Epistemic Uncertainty
Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty. Experts $B_{1...N}A_{1...N}$ capture functional relationships in the data throughout fine-tuning, the weights $w_{1...N}|h$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$. In the UQ4CT workflow, we align this uncertainty with predictive correctness. When the router makes a correct prediction, the loss reinforces this decision, thereby increasing the router confidence in its selection, which aligns with a lower epistemic uncertainty. Conversely, when the router makes an incorrect prediction, the loss penalizes this selection, potentially causing the router to distribute its probabilities more broadly across experts, which is indicative of higher epistemic uncertainty.
Motivated by the decomposition in Eq. (3) for least squares loss, we may consider a general distance $d$, and define epistemic uncertainty that characterizes the variation caused by model training procedure. Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs. Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(M|x)}\mathbb{E}_{M^{\prime}\sim P(M^{\prime}|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right].$ (5)
3.2 Quantifying FEU with LoRA MoE Framework
We quantify the functional-level epistemic uncertainty (FEU) with the MoE architecture, which is represented by the embedding $e(a)$ in Eq. (5). As shown in Figure 1, the LoRA experts $B_{1...N}A_{1...N}$ capture important functional relationships in the data during fine-tuning. We treat these functions represented by the LoRA experts as basis functions. Conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions. Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions. In particular, the weights $w_{1...N}|h$ from the top-k router dynamically selects these basis functions conditioned on the input hidden state $h$. The top-k weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection.
In this mixture of LoRA experts architecture, we follow the routing mechanisms of the MoE layers as in Eq. (2) and (3). Specifically, we employ top-2 gate routers, which chooses the $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (6)
Given an input prompt $x$ with length $s$, we model functional-level epistemic uncertainty (FEU) by aggregating $R^{\ell}(\mathbf{h}^{\ell}_{i})$ over both layer dimension and sequence dimension:
$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right]$ (7)
3.3 Training Calibration Loss
We then calibrate the FEU model of the epistemic uncertainty against predictive accuracy. Specifically for the MoE top-k routers, we design the following calibration loss for training:
$\mathcal{L}_{\mathrm{cal}}=\left\|\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FEU}(x)\right\|^{2},$ (8)
where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$. As shown in Figure 1, this term effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness. Therefore, this calibrated training loss ensures that the router’s certainty aligns more closely with the actual confidence the model should have given the available data and its knowledge state. Thus, it effectively calibrates the router’s decisions in line with epistemic uncertainty.
Load balancing is a common technique to ensure even exploitation across experts with the MoE architecture (Fedus et al., 2022). We follow the load balancing loss $\mathcal{L}_{b}$ proposed by (Li et al., 2024) and define our loss function as:
$\mathcal{L}=\mathbf{CE}+\alpha\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (9)
where $\mathbf{CE}$ represents cross entropy loss, $\alpha$ and $\beta$ are the hyperparameters of two auxiliary terms.
4 Related Work
4.1 Parameter-Efficient Fine Tuning for LLMs
Large Language Models (Brown et al., 2020; Chowdhery et al., 2022; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks. Building on these advances, instruction fine-tuning(Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems(Wu et al., 2023b; Achiam et al., 2023b). However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT)(Mangrulkar et al., 2022), model distillation(Liu et al., 2023; Xiao et al., 2023), quantization(Frantar et al., 2022; Xiao et al., 2022b), and pruning(Frantar and Alistarh, 2023; Ma et al., 2023). Among these, LoRA(Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference. For example, VeRA(Kopiczko et al., 2023) introduces learnable scaling vectors to modify shared pairs of frozen random matrices across layers, while FedPara(Hyeon-Woo et al., 2021) focuses on low-rank Hadamard products for federated learning settings. Tied-Lora(Renduchintala et al., 2023) applies weight tying to further minimize the number of trainable parameters. AdaLoRA(Zhang et al., 2023) uses Singular Value Decomposition (SVD) to prune less important singular values for efficient updates, and DoRA(Liu et al., 2024) separates pre-trained weights into magnitude and direction components, applying LoRA to update the directional component during fine-tuning, thus reducing the number of parameters to be trained.
4.2 Mixture of Experts
The Mixture-of-Experts concept(Jacobs et al., 1991), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data. Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
Different MoE architectures have since emerged, distinguished by their expert sampling and routing strategies. For example, LLaVA-MoLE(Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines. Other MoE-based approaches include MoRAL(Yang et al., 2024b), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE(Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting. PESC(Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational and GPU memory requirements. MoE-LoRA(Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE(Gou et al., 2023) activates task-specific model parameters based on instruction clusters. MixLoRA (Li et al., 2024) implements a high-throughput framework for LoRA MoE training and inference process, constructing LoRAs as stochastic experts to reduce computational overhead while expanding model capacity.
Despite the performance improvements these architecture advancements have brought, the overconfidence problem of fine-tuned models is lacking attention (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Enhancing the uncertainty estimation capabilities of these models is fundamental toward more reliable, interpretable and trustworthy applications of LLMs.
4.3 Uncertainty Quantification in LLMs
Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks(Gal and Ghahramani, 2015, 2016b; Malinin and Gales, 2018; Ovadia et al., 2019; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023). This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research(Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023). LLMs, particularly in generative tasks, pose unique challenges, especially when it comes to measuring the uncertainty of their outputs(Liu et al., 2019; Malinin and Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs(Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
Existing works have investigated the application of ensembling in fine-tuning LLMs for uncertainty quantification. Gleave and Irving (2022); Sun et al. (2022) focus on uncertainty estimation in full model fine-tuning, while this approach inherently incurs significant memory overhead. Wang et al. (2023); Zhai et al. (2023b); Balabanov and Linander (2024) explore the use of LoRA ensembles for uncertainty estimation in LLMs. Yang et al. (2024a) applies a post-hoc Laplace approximation Mackay (1992) to model LoRA parameters after fine-tuning. BatchEnsemble(Wen et al., 2020), introduces component-specific rank-1 matrices as multiplicative modifications to a base model. Though this method has been applied to LLMs, it has been used in the pre-training phase rather than fine-tuning(Tran et al., 2022). None of these methods provide calibrations on epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
5 Experiments
5.1 Datasets
We include $5$ multiple-choice question answering benchmarks to evaluate UQ4CT: OpenBookQA (OBQA, Mihaylov et al. (2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science. We also use computer science, law, medication and engineering subsets from MMLU dataset(Hendrycks et al., 2020) to evaluate performance under distribution shift. We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance.
Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across $4$ common sense reasoning tasks and a domain-specific task. UQ4CT shows significant ECE improvements while maintaining high accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ LoRA $69.5_{1.93}$ $74.8_{1.39}$ $53.8_{0.6}$ $72.1_{0.87}$ $59.9_{2.13}$
MC Drop $66.8_{3.66}$ $76.8_{1.30}$ $50.9_{2.01}$ $74.8_{1.34}$ $58.2_{2.11}$
Ensemble $66.2_{3.7}$ $71.2_{1.0}$ $47.5_{0.57}$ $75.5_{1.4}$ $59.6_{6.9}$
LA $68.7_{1.32}$ $74.6_{2.11}$ $51.4_{0.83}$ $70.8_{1.24}$ $55.2_{3.29}$
MixLoRA $71.5_{1.05}$ $77.7_{2.27}$ $54.3_{1.07}$ $75.5_{2.91}$ $61.6_{1.76}$
UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
ECE $\downarrow$ LoRA $11.9_{0.78}$ $11.9_{2.04}$ $19.4_{4.75}$ $10.2_{1.07}$ $14.3_{0.64}$
MC Drop $12.2_{0.85}$ $11.9_{1.99}$ $19.8_{4.85}$ $10.9_{0.24}$ $14.3_{0.56}$
Ensemble $7.28_{2.3}$ $9.1_{1.49}$ $10.23_{1.39}$ $8.83_{2.35}$ $13.5_{3.29}$
LA $17.1_{1.72}$ $16.6_{3.7}$ $18.1_{0.5}$ $17.2_{1.2}$ $12.6_{1.9}$
MixLoRA $7.88_{2.09}$ $9.09_{0.81}$ $10.74_{1.07}$ $12.9_{1.99}$ $12.5_{1.32}$
UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
5.2 Experiment Setup
We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024) and compare the average performance in three random runs and report the mean and standard deviation with following baselines. We use the publicly available LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model. In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-fowrward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer). Details are provided in Appendix A.1
  • LoRA(Hu et al., 2021a). We use standard LoRA fine-tuning as lower performance bound.
  • Monte Carlo (MC) Dropout(Gal and Ghahramani, 2016a) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
  • Deep Ensemble(Lakshminarayanan et al., 2017) averages the predictions from each ensemble member which have been trained with varying random initialization. We combine deep ensemble with LoRA by fine-tuning $3$ randomly initialized LoRAs together and ensembling their output as final predictions.
  • Laplace-LoRA (LA)(Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
  • MixLoRA(Li et al., 2024) incorporates LoRAs via a router network to reduce computational overhead while expanding model capacity. We add this as a baseline to resemble plain LoRA MoE model performance.
Evaluation Metrics.
We measure the prediction accuracy on the validation set for all $5$ tasks. For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al. (2017)) with bin size of $15$, which measures the alignment between predicted probabilities and empirical accuracy. We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications. Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts. Metric details are provided in Appendix A.2.
5.3 Results
We assess the prediction accuracy and uncertainty calibration of models under both in-distribution and distribution shift scenarios. The in-distribution scenario examines the alignment of the fine-tuned model on the target downstream task, while the distribution shift scenario evaluates the generalizability of the model on novel tasks beyond the fine-tuned domain. These two scenarios combined provides a comprehensive assessment of model robustness in real-world applications where it is essential for the model to excel on its primary task while maintaining the ability to effectively handle unforeseen or out-of-distribution inputs.
Table 2: Performance comparison of different methods fine-tuned on OBQA dataset with LlaMA2-7B across $2$ smaller distribution shift (DS) tasks and $4$ larger distribution shift tasks. UQ4CT shows significant ECE improvements while maintaining high accuracy.
ID Smaller DS Larger DS
Metrics Methods OBQA ARC-C ARC-E CS Eng Law Health
ACC $\uparrow$ LoRA $72.1_{0.87}$ $58.6_{1.93}$ $66.5_{3.38}$ $35.5_{2.35}$ $30.8_{1.72}$ $34.9_{1.41}$ $39.1_{1.52}$
MC Drop $74.8_{1.34}$ $58.7_{2.07}$ $66.6_{3.30}$ $36.0_{1.69}$ $30.3_{2.25}$ $35.1_{0.86}$ $39.1_{1.35}$
Ensemble $75.5_{1.4}$ $57.7_{0.78}$ $69.1_{0.48}$ $36.7_{2.18}$ $30.3_{1.13}$ $35.3_{1.02}$ $39.9_{1.89}$
LA $70.8_{1.24}$ $58.7_{0.58}$ $67.9_{0.41}$ $33.7_{1.22}$ $29.6_{1.32}$ $35.4_{0.75}$ $38.5_{1.61}$
MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $69.2_{1.02}$ $35.2_{2.92}$ $30.3_{0.98}$ $35.9_{0.43}$ $40.6_{1.13}$
UQ4CT $77.3_{1.36}$ $58.8_{1.06}$ $65.8_{1.31}$ $36.2_{1.24}$ $34.1_{2.31}$ $35.8_{1.01}$ $40.0_{1.24}$
ECE $\downarrow$ LoRA $10.2_{1.07}$ $16.7_{2.28}$ $13.3_{2.48}$ $29.7_{2.69}$ $32.3_{1.85}$ $29.2_{3.08}$ $31.0_{2.13}$
MC Drop $10.9_{0.24}$ $16.7_{2.20}$ $13.2_{2.21}$ $23.2_{2.32}$ $31.6_{1.64}$ $28.0_{2.93}$ $25.9_{2.27}$
Ensemble $8.83_{2.35}$ $15.1_{1.09}$ $11.1_{0.99}$ $22.4_{1.32}$ $28.5_{2.13}$ $29.0_{1.37}$ $24.5_{0.39}$
LA $17.2_{1.2}$ $16.2_{0.5}$ $24.4_{0.42}$ $28.6_{2.61}$ $30.5_{1.43}$ $29.5_{1.83}$ $30.7_{1.2}$
MixLoRA $12.9_{1.99}$ $19.0_{1.88}$ $14.5_{2.57}$ $26.4_{3.25}$ $33.7_{1.87}$ $30.3_{2.27}$ $28.3_{1.07}$
UQ4CT $\mathbf{5.0_{1.15}}$ $\mathbf{8.9_{3.46}}$ $\mathbf{6.5_{1.85}}$ $\mathbf{19.6_{2.90}}$ $\mathbf{23.1_{1.17}}$ $\mathbf{25.9_{3.43}}$ $\mathbf{21.9_{3.49}}$
5.3.1 In-distribution Performance
We first evaluate UQ4CT and baseline models fine-tuned on the $4$ common sense reasoning tasks and the climate question answering task under the in-distribution scenario, where models are trained and evaluated on different splits of the same dataset. Note that one of the key advantages of UQ4CT is that uncertainty calibration happens during the fine-tuning stage with little computational overhead, while other UQ methods require repetitive sampling or other post-hoc complexities.
As shown in Table 1, UQ4CT demonstrates notable improvements in uncertainty calibration across a variety of tasks. Across all benchmarks, UQ4CT maintains competitive accuracy (ACC) compared to the baseline methods. For example, on the BoolQ and ClimateQA datasets, UQ4CT achieves accuracy rates of $73.5\%$ and $63.3\%$, respectively. This empirically demonstrates that UQ4CT is capable of maintaining high accuracy with uncertainty calibration, which assures the gain in UQ performance does not compromise accuracy.
The most significant performance improvement is observed in the reduction of Expected Calibration Error (ECE). UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks. Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration. This improvement is critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
5.3.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios. Similar to the dataset setup in (Yang et al., 2024a), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA. For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health. These tasks have very broad coverage of the domain task at various knowledge levels ranging from elementary school to professionals. This domain-specificity demonstrates larger distribution shift from OBQA, which is a general common sense reasoning task. Details of the ensemble is provided in Appendix. A.3.
The distribution shift evaluations are provided in Table 2. UQ4CT provides significant improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios. For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario. For the more challenging larger distribution shifts, UQ4CT still achieves the best ECE performance among all baseline models. Note that UQ4CT also achieves competitive prediction accuracy across all domain-specific tasks. This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
5.4 Ablation Study
5.4.1 Incremental Weighting on Calibration Term
Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data. Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration. We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
$\beta=\min\left\{1,\frac{\mathrm{current\_grad\_step}}{50}\right\}.$ (10)
We choose $50$ gradient steps from our observation that training loss generally stabilizes after $50$ gradient steps, indicating the LoRA experts have learned some functional relationships from data.
As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks. This demonstrates the advantage of uncertainty calibration even in the early stage. In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty. Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintains similar accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
6 Discussion & Conclusion
In this work, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which addresses the overconfidence issues commonly encountered during fine-tuning of large language models. We present a functional perspective on quantifying epistemic uncertainty in LLMs and utilize it for uncertainty-calibrated fine-tuning. By incorporating functional-level epistemic uncertainty quantification with a mixture-of-experts framework, our proposed uncertainty-calibrated training loss effectively addresses the challenge of overconfidence in fine-tuned LLMs by significantly improving uncertainty calibration while maintaining high accuracy. Our evaluations demonstrate that UQ4CT reduces the Expected Calibration Error by more than 25% without compromising accuracy across a variety of downstream tasks, including common-sense and domain-specific reasoning, under in-distribution and out-of-distribution scenarios.
The limitation of UQ4CT lies in its dependency on predictive correctness. For general language modeling tasks such as chat completion, there lacks a clear metric on whether the response is correct or not. This limits the application of UQ4CT as naively token matching is a poor indicator of semantic correctness due to the ambiguous nature of language. For future work, we plan to continue exploring ways to adapt UQ4CT on open-ended problems that lacks a definitive metric.
Reproducibility Statement
We share our experimental details in Appendix A.1, and also provide the code and model weights for running experiments in the supplementary materials to reproduce our model performance results.
Acknowledgement
This work was supported in part by U. S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, and NSF Grants #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069, DARPA AIE FoundSci and DARPA YFA.
Appendix A Appendix
A.1 Training Details
We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability. For each expert, we use $rank=16$ and $alpha=32$. We use batch size of $16$ to train our model. For climate task, we set the learning rate to $45e-$ 4 and dropout rate to $0.1$ to incorporate the small dataset size. For other tasks, we use $42e-$ 4 as our learning rate with dropout $0.05$. We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training.
The experimental setup for single LoRA based models is similar with LoRA ranks set to $80$ to accommodate the MoE model size. For the ensemble baseline, we use an ensemble size of $8$ with $rank=16$. For Laplace-LoRA, we follow the Laplace hyperparameters in this Github Repository.
A.2 Expected Calibration Error
Expected calibration error (ECE) is a commonly used metric to asses uncertainty quantification performance. ECE measures the alignment between prediction accuracy and model confidence through regrouping the predicted probabilities into $m$ bins. This method then computes the weighted average of the difference between average accuracy and confidence in each bin:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\text{acc}(B_{m})-\text{conf}(B_{m})|,$ (11)
where $|B_{m}|$ is the number of evaluated datapoints in bin $m$, acc and conf is calculated as following:
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i}),$ (12)
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}P(\hat{y}_{i}).$ (13)
A.3 MMLU Distribution Shift Dataset Composition
  • Computer Science (CS):
    • College Computer Science
    • Computer Security
    • High School Computer Science
    • Machine Learning
  • Engineering (Eng):
    • Electrical Engineering
  • Law:
    • International Law
    • Jurisprudence
    • Professional Law
  • Health:
    • Anatomy
    • Clinical Knowledge
    • College Medicine
    • Human Aging
    • Nutrition
    • Professional Medicine
    • Virology

### v2

@@
   1.
-From common-sense reasoning to domain-specific tasks, parameter-efficient fine tuning (PEFT) methods for large language models (LLMs) have showcased significant performance improvements on downstream tasks.
-However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions, particularly due to sparse training data.
-This overconfidence reflects poor epistemic uncertainty calibration, which arises from limitations in the model’s ability to generalize with limited data.
-Existing PEFT uncertainty quantification methods for LLMs focus on the post fine-tuning stage and thus have limited capability in calibrating epistemic uncertainty.
-To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates functional-level epistemic uncertainty during the fine-tuning stage via a mixture-of-expert framework.
-We show that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks.
-Furthermore, UQ4CT maintains superior ECE performance with high accuracy under distribution shift, showcasing improved generalizability.
+Accurate uncertainty quantification of large language models (LLMs) provides credibility measure over their outputs.
+However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions due to the limitations in the models’ ability to generalize with limited data.
+Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage and fall short of calibrating epistemic uncertainty.
+To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates epistemic uncertainty over the space of functions that map input prompts to outputs.
+We implement UQ4CT during the fine-tuning stage via a mixture-of-experts framework that hierarchically decomposes the functional space.
+We demonstrate that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks.
+Even under distribution shift, UQ4CT maintains superior ECE performance with high accuracy, showcasing improved generalizability.
 11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint.
@@
 1 Introduction
-Large Language Models (LLMs) have revolutionized various domains as general task solvers (Chang et al., 2024).
-To adapt LLMs for specific downstream tasks or create instruction-following models, fine-tuning have become increasingly important (Houlsby et al., 2019; Hu et al., 2021a; Liu et al., 2022; Ding et al., 2022, 2023).
-This involves additional training on pre-trained LLMs using a smaller dataset (Zhong et al., 2021; Ren et al., 2022).
-Through fine-tuning, the model parameters are updated to better adapt to the domain-specific knowledge (Peng et al., 2023).
-To reduce the computational cost for fine-tuning, Hu et al.
-(2021a) proposed Low-Rank Adaptation (LoRA), which effectively reduces the parameters required for fine-tuning by introducing low-rank trainable matrices at each layer of the transformer architecture instead of fine-tuning the full model parameters.
-Li et al.
-(2024); Wu et al.
-(2024a) proposed LoRA Mixture-of-Experts (MoE) models which grants better performance while maintaining parameter efficiency.
-However, previous studies have shown that fine-tuned LLMs are often overconfident with their predictions (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023).
-This resembles poorly calibrated uncertainty (Zhou et al., 2022) due to the sparsity of fine-tuning data.
-Overconfidence is a crucial problem in safety-related decision making or in fields where data is very limited, such as experimental design, climate science and medical diagnosis (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022).
-Thus, methods that enhance uncertainty quantification of fine-tuned LLMs is urgently needed to assure trustworthy predictions for better application.
-Established uncertainty quantification methods have been studied in conjunction with the LoRA structure.
-Monte-Carlo dropout (Gal and Ghahramani, 2016a) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them.
-Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty.
-Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023a) combines the predictions from multiple LoRA adapters for better uncertainty calibration.
-Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
-Although these methods have demonstrated improved uncertainty estimations, they all utilize a fixed set of LoRA parameters fine-tuned over the entire downstream task dataset.
-The point estimates of parameters have very limited capabilities in capturing epistemic uncertainty, while direct calibration of epistemic uncertainty over the entire LoRA parameter space is an ideal but not practical approach.
-Therefore, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the functional-level epistemic uncertainty via the LoRA MoE architecture.
-We propose a functional perspective on LoRA parameters where we treat the LoRA experts as basis functions and consider the more complex, prompt dependent functions as mixtures of those basis functions.
-On top of learning the parameters in the LoRA experts, UQ4CT also trains a prompt-dependent LoRA mixture to form a calibrated distribution over the functional space.
-The LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input.
+Quantifying credibility of the outputs has been one of the most important problems around large language models (LLMs) (Chang et al., 2024).
+In particular, fine-tuned LLMs often struggle with overconfidence in gauging the uncertain of their outputs—oftentimes due to limited training data—and does not reflect the true credibility of the answers that they generate (Xiao et al., 2022a; Zhou et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023).
+Such overconfidence has become a crucial problem in safety-related decision making and in fields where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022).
+Methods that enhance uncertainty quantification of fine-tuned LLMs are sought after for the assurance of trustworthy predictions in applications.
+One salient challenge of uncertainty quantification in large models is the trade-off among accuracy, calibration, and tractability.
+One would first like to maintain the capability of the LLMs and not sacrifice accuracy of the generated outputs.
+Given this desideratum, a number of existing works focus on perturbing the prompts that are fed into the language models and quantify the uncertainty associated with such perturbations (Hou et al., 2023; Farquhar et al., 2024; Gao et al., 2024).
+These methods capture the aleatoric uncertainty assuming that the trained model corresponds to the ground-truth data distribution.
+However, these approaches cannot capture the epistemic uncertainty that arises from the model being trained on limited data.
+To capture this epistemic uncertainty, one would need to introduce changes or perturbations to the model in accordance with the data distribution.
+On the tractability side, it is not realistic to consider perturbations to all the model parameters for an LLM (Minaee et al., 2024).
+We therefore propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the epistemic uncertainty in LLMs.
+Since we only focus on the epistemic uncertainty over the functional space that maps input prompts to the outputs, we no longer need to consider uncertainty in the entire parameter space.
+Instead, we only need to measure the uncertainty regarding the functional relationships.
+Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the interesting functional relationships arising from the underlying data.
+In particular, we make use of the low rank adaptation (LoRA, Hu et al.
+(2021a)) ensemble at each layer of the LLM to form the basis functions.
+We then utilize the mixture of experts (MoE) architecture (Li et al., 2024) to hierarchically combine the basis functions and form the functional space.
+During fine-tuning, UQ4CT simultaneously learns the parameters in the LoRA experts and the prompt-dependent mixture weights to output the calibrated distribution over the functional space.
+During inference, the LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input.
 The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space.
 We calibrate functional level epistemic uncertainty to align with predictive correctness during training time.
-This significantly improves uncertainty estimations of the model on its predictions without compromising the accuracy.
+This approach significantly improves uncertainty estimations of the model on its predictions without compromising accuracy.
 To summarize, our contributions include:
-  • A novel epistemic uncertainty quantification approach with Mixture-of-Experts architecture during fine-tuning stage to model functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
-  • A novel training calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
-  • More than $25\%$ Expected Calibration Error reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task, superior ECE performance under distribution shift scenarios on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
+  • A novel epistemic uncertainty quantification approach with MoE architecture during fine-tuning to quantify functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
+  • A novel calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
+  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance under distribution shifts on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
 2 Preliminaries
@@
 2.2 Mixture of Experts (MoE)
-LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is a efficient approach to scale the number of parameters while maintaining the same computational bounds.
+LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is an efficient approach to scale the number of parameters while maintaining the same computational bounds.
 LoRA MoE utilizes the top-k router to assign each token to the LoRA experts (Lepikhin et al., 2020).
 The router is a linear layer that maps the input hidden state $\mathbf{h}$ to a probability distribution of candidate experts.
+The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and feed forward layer (FFN), together with residual connections.
+Formally, given $\mathbf{h}_{o}=[s_{0},s_{1},...,s_{n}]$ as the tokenized input text, the output of $\ell$-th layer is generated as:
+$\mathbf{z}^{\ell}=f_{attn}(\text{FFN}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{FFN}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (2)
+Here, $f_{attn}$ represents the attention calculation step.
 Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension.
-With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router chooses $k$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
-$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-k}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (2)
+With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-n gate router $R(\cdot)$ chooses n experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
+$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-n}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (3)
 Finally, we obtain the final MixLoRA prediction with:
-$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{i}\mathbf{A}^{\ell}_{i}\cdot\mathbf{h}^{\ell}$ (3)
-where $\mathbf{W}$ is the pretrained weights of the feed-forward network (FFN) layer and $\mathbf{B}^{\ell}_{i}\mathbf{A}^{\ell}_{i}$ is the i-th LoRA expert.
+$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
+where $\mathbf{W}$ is the pretrained weights of the target fine-tuning layer and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
+With MixLoRA defined in Equation 4, we can apply MixLoRA layers at $q,k,v$ encoding and FFN layers:
+$\mathbf{z}^{\ell}=f_{attn}(\text{MixLoRA}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{MixLoRA}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (5)
+Note that the calculation of $E^{\ell}_{k}$ involves $\mathbf{W}^{\ell}$, which represents the weight of the original ATTN or FFN layer.
 2.3 Aleatoric and Epistemic Uncertainties
@@
 However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a).
-A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target downstream task (Lin et al., 2023; Luo et al., 2023).
+A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target task (Lin et al., 2023; Luo et al., 2023).
 Motivated by these observations, we explore functional-level epistemic uncertainty in generative models and aim to develop metrics that assess model performance on specific problem instances to fine-tune the parameter mixture.
@@
 Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$.
-We hence name the task “functional-level uncertainty quatification”.
+We hence name the task “functional-level uncertainty quantification”.
 3.1 Functional-level Epistemic Uncertainty
 Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty.
-Experts $B_{1...N}A_{1...N}$ capture functional relationships in the data throughout fine-tuning, the weights $w_{1...N}|h$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$.
+Experts $B_{1...K}A_{1...K}$ capture functional relationships in the data throughout fine-tuning, the weights $\alpha_{1...K}$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$.
 In the UQ4CT workflow, we align this uncertainty with predictive correctness.
@@
 Motivated by the decomposition in Eq.
-(3) for least squares loss, we may consider a general distance $d$, and define epistemic uncertainty that characterizes the variation caused by model training procedure.
+(3) for least squares loss, we consider a general distance $d$, and formally define epistemic uncertainty as the variance caused by model training procedure.
 Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs.
 Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
-$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(M|x)}\mathbb{E}_{M^{\prime}\sim P(M^{\prime}|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right].$ (5)
+$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(\cdot|x)}\mathbb{E}_{M^{\prime}\sim P(\cdot|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right]$ (7)
+Here, $a^{\prime}$ and $a$ represents the outputs sampled from the parameter mixtures $M^{\prime}$ and $M$ conditioned on the prompt $x$.
+The epistemic uncertainty measures the least squares distance between $e(a)$ and $e(a^{\prime})$.
+3.2 Decomposition of the Functional Space
+Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the functional relationships in the underlying data.
+Then the functional space can be decomposed into a mixture of the LoRA ensembles.
+Naively, this decomposition writes:
+$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{1}=1}^{K}\sum_{k^{2}=1}^{K}\cdots\sum_{k^{L}=1}^{K}\alpha_{k^{1},k^{2},\dots,k^{L}}g_{k^{L}}^{L}\left(g_{k^{L-1}}^{L-1}\left(\dots g_{k^{1}}^{1}(\mathbf{x})\right)\right),$ (8)
+$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right),$ (9)
+where $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e.
+layer norm, attention calculation, etc.).
+This linear decomposition requires examining $K^{L}$ parameters to span the entire functional space.
+We instead consider a hierarchical decomposition of $f(\mathbf{x})$:
+$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\Bigg{(}\sum_{k^{L-1}=1}^{K}\alpha^{L-1}_{k^{L-1}}g_{k^{L-1}}^{L-1}\Bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$
+Here the output function is determined by the combinations of the weights $\alpha^{\ell}_{k^{\ell}}$ across all the layers.
+As a result, we only need $K\cdot L$ parameters to represent various compositions of the functional relationships.
+To further improve the efficiency of representation, we apply a sparse dynamic routing and choose the mixture weights $\alpha_{k^{\ell}}=R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}$ that depend on the input from the previous layer:
+$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
 3.3 Quantifying FEU with LoRA MoE Framework
@@
 (7).
-As shown in Figure 1, the LoRA experts $B_{1...N}A_{1...N}$ capture important functional relationships in the data during fine-tuning.
-We treat these functions represented by the LoRA experts as basis functions.
-Conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions.
+As shown in Figure 1, the LoRA experts that are made of matrices $B^{1...L}_{1...K},A^{1...L}_{1...K}$ in Eq.
+(1) capture important functional relationships in the data during fine-tuning.
+We treat these functional relationships represented by the LoRA experts as basis functions.
+As shown in Eq.
+(10), conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions, hierarchically through each layer.
 Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions.
-In particular, the weights $w_{1...N}|h$ from the top-k router dynamically selects these basis functions conditioned on the input hidden state $h$.
-The top-k weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection.
-In this mixture of LoRA experts architecture, we follow the routing mechanisms of the MoE layers as in Eq.
+As shown in Figure 1, the weights $\alpha_{k}^{\ell}$ from the top-n router dynamically mix these basis functions given the input hidden state.
+The top-n weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection.
+With mixture of LoRA experts architecture, we follow routing mechanisms of the MoE layers as in Eq.
 (3) and (4).
-Specifically, we employ top-2 gate routers, which chooses the $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
-$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (6)
+Specifically, we employ top-2 gate routers, which chooses $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
+$\vspace{-0.5mm}R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (11)
 Given an input prompt $x$ with length $s$, we model functional-level epistemic uncertainty (FEU) by aggregating $R^{\ell}(\mathbf{h}^{\ell}_{i})$ over both layer dimension and sequence dimension:
-$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right]$ (7)
+$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right].$ (12)
 3.4 Training Calibration Loss
-We then calibrate the FEU model of the epistemic uncertainty against predictive accuracy. Specifically for the MoE top-k routers, we design the following calibration loss for training:
+The FEU model of the epistemic uncertainty enables us to calibrate the parameter mixtures against predictive accuracy for better aligned output distributions.
+Specifically for the MoE top-k routers, we design the following calibration loss for training:
 $\mathcal{L}_{\mathrm{cal}}=\left\|\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FEU}(x)\right\|^{2},$ (13)
 where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$.
-As shown in Figure 1, this term effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness.
-Therefore, this calibrated training loss ensures that the router’s certainty aligns more closely with the actual confidence the model should have given the available data and its knowledge state.
-Thus, it effectively calibrates the router’s decisions in line with epistemic uncertainty.
-Load balancing is a common technique to ensure even exploitation across experts with the MoE architecture (Fedus et al., 2022).
-We follow the load balancing loss $\mathcal{L}_{b}$ proposed by (Li et al., 2024) and define our loss function as:
-$\mathcal{L}=\mathbf{CE}+\alpha\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (9)
-where $\mathbf{CE}$ represents cross entropy loss, $\alpha$ and $\beta$ are the hyperparameters of two auxiliary terms.
+We employ a one-hot definition of the ground truth confidence.
+When the prediction from current mixture model matches the ground truth, the ground truth confidence is $1$.
+Otherwise, when the predictions do not match, the ground truth confidence is 00.
+3.5 Calibrating the Mixture
+Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration with functional-level epistemic uncertainty.
+As shown in Figure 1, $\mathcal{L}_{\mathrm{cal}}$ effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness.
+Ideally, when the $K$ LoRA experts together capture all the functional relationships across the data distribution with cross entropy during fine-tuning, our proposed loss $\mathcal{L}_{cal}$ also finds proper mixture of LoRA experts conditioned on the input $x$ by conditionally promoting expert exploitation and exploration.
+This allows the model to select correct functional relationships regarding $x$ to generate an output that better matches the data distribution, which grants calibrated uncertainty estimations.
+Load balancing is a common regularization to ensure even utilization across experts with the MoE architecture (Fedus et al., 2022).
+We follow the load balancing loss $\mathcal{L}_{b}$ proposed by Li et al.
+(2024) and define our loss function as:
+$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
+where $\mathbf{CE}$ represents cross entropy loss, $\gamma$ and $\beta$ are the hyperparameters of two auxiliary terms. Details about $\mathcal{L}_{b}$ can be found in Appendix A.1.
 4 Related Work
 4.1 Parameter-Efficient Fine Tuning for LLMs
-Large Language Models (Brown et al., 2020; Chowdhery et al., 2022; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks.
+Large Language Models (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks.
 Building on these advances, instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems (Wu et al., 2023b; Achiam et al., 2023b).
 However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
-To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT)(Mangrulkar et al., 2022), model distillation(Liu et al., 2023; Xiao et al., 2023), quantization(Frantar et al., 2022; Xiao et al., 2022b), and pruning(Frantar and Alistarh, 2023; Ma et al., 2023).
+To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT) (Mangrulkar et al., 2022), model distillation (Liu et al., 2023; Xiao et al., 2023), quantization (Frantar et al., 2022; Xiao et al., 2022b), and pruning (Ma et al., 2023).
 Among these, LoRA (Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference.
@@
 4.2 Mixture of Experts
-The Mixture-of-Experts concept(Jacobs et al., 1991), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data.
+The Mixture-of-Experts concept (Jacobs et al., 1991; Wang et al., 2020), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data.
 Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
@@
 For example, LLaVA-MoLE (Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines.
-Other MoE-based approaches include MoRAL(Yang et al., 2024b), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE(Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting.
-PESC(Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational and GPU memory requirements.
+Other MoE-based approaches include MoRAL (Yang et al., 2024a), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE (Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting.
+PESC (Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational costs.
 MoE-LoRA (Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE (Gou et al., 2023) activates task-specific model parameters based on instruction clusters.
@@
 4.3 Uncertainty Quantification in LLMs
-Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks(Gal and Ghahramani, 2015, 2016b; Malinin and Gales, 2018; Ovadia et al., 2019; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023).
+Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks (Gal and Ghahramani, 2015, 2016a; Malinin and Gales, 2018; Ovadia et al., 2019; Dusenberry et al., 2020; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023).
 This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research (Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023).
@@
 The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs (Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
-Existing works have investigated the application of ensembling in fine-tuning LLMs for uncertainty quantification.
-Gleave and Irving (2022); Sun et al.
-(2022) focus on uncertainty estimation in full model fine-tuning, while this approach inherently incurs significant memory overhead.
-Wang et al.
-(2023); Zhai et al.
-(2023b); Balabanov and Linander (2024) explore the use of LoRA ensembles for uncertainty estimation in LLMs.
-Yang et al.
-(2024a) applies a post-hoc Laplace approximation Mackay (1992) to model LoRA parameters after fine-tuning.
-BatchEnsemble(Wen et al., 2020), introduces component-specific rank-1 matrices as multiplicative modifications to a base model.
-Though this method has been applied to LLMs, it has been used in the pre-training phase rather than fine-tuning(Tran et al., 2022).
-None of these methods provide calibrations on epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
+Established uncertainty quantification methods have been studied in conjunction with the LoRA structure.
+Monte-Carlo dropout (Gal and Ghahramani, 2016b) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them.
+Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty.
+Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration.
+Laplace-LoRA (Yang et al., 2024b) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
+However, none of these methods provide calibrations regarding epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
 5 Experiments
 5.1 Datasets
-We include $5$ multiple-choice question answering benchmarks to evaluate UQ4CT: OpenBookQA (OBQA, Mihaylov et al.
+We include $5$ multiple-choice question answering benchmarks: OpenBookQA (OBQA, Mihaylov et al.
 (2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science.
@@
 We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance.
-Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across $4$ common sense reasoning tasks and a domain-specific task.
-UQ4CT shows significant ECE improvements while maintaining high accuracy.
+We report the average model performance over $3$ random runs and the standard deviations in the subscript.
+Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task.
+UQ4CT shows substantial ECE improvements while maintaining high accuracy.
 Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
@@
   • LoRA (Hu et al., 2021a). We use standard LoRA fine-tuning as lower performance bound.
-  • Monte Carlo (MC) Dropout(Gal and Ghahramani, 2016a) keeps dropout on at both training and testing time.
+  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016b) keeps dropout on at both training and testing time.
 By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions.
@@
 We combine deep ensemble with LoRA by fine-tuning $3$ randomly initialized LoRAs together and ensembling their output as final predictions.
-  • Laplace-LoRA (LA)(Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
+  • Laplace-LoRA (LA) (Yang et al., 2024b) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
   • MixLoRA (Li et al., 2024) incorporates LoRAs via a router network to reduce computational overhead while expanding model capacity.
@@
 For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al.
-(2017)) with bin size of $15$, which measures the alignment between predicted probabilities and empirical accuracy.
+(2017)) with $15$ bins, which measures the alignment between predicted probabilities and empirical accuracy.
 We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications.
 Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts.
-Metric details are provided in Appendix A.2.
+Metric details are given in Appendix A.6.
 5.3 Results
@@
 These two scenarios combined provides a comprehensive assessment of model robustness in real-world applications where it is essential for the model to excel on its primary task while maintaining the ability to effectively handle unforeseen or out-of-distribution inputs.
-Table 2: Performance comparison of different methods fine-tuned on OBQA dataset with LlaMA2-7B across $2$ smaller distribution shift (DS) tasks and $4$ larger distribution shift tasks.
-UQ4CT shows significant ECE improvements while maintaining high accuracy.
+Table 2: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA2-7B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks.
+UQ4CT shows substantial ECE improvements while maintaining high accuracy.
 ID Smaller DS Larger DS
@@
 This empirically demonstrates that UQ4CT is capable of maintaining high accuracy with uncertainty calibration, which assures the gain in UQ performance does not compromise accuracy.
-The most significant performance improvement is observed in the reduction of Expected Calibration Error (ECE).
+The most substantial performance improvement is observed in the reduction of Expected Calibration Error (ECE).
 UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks.
 Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration.
-This improvement is critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
+In addition to experiments on LLaMA-2-7B in the main text, we also present additional experiments on fine-tuning Mistral-7B in Appendix A.2 for more comprehensive evaluation of our method.
+For both LLaMA-2-7B and Mistral-7B models, UQ4CT consistently shows substantial improvements in uncertainty calibration.
+The improvements are critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
 5.3.2 Performance Under Distribution Shift
@@
 Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios.
-Similar to the dataset setup in (Yang et al., 2024a), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA.
+Similar to the dataset setup in (Yang et al., 2024b), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA.
 For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health.
@@
 Details of the ensemble is provided in Appendix.
-A.3.
+A.7.
 The distribution shift evaluations are provided in Table 2.
-UQ4CT provides significant improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios.
+UQ4CT provides substantial improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios.
 For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario.
@@
 This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
+Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintaining similar accuracy.
+Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
+ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
+UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
+ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
+UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
 5.4 Ablation Study
+In this section, we conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$.
+We first evaluate the incremental weighting performance of the calibration term, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning.
+Then, we perform a sensitivity test, where we explore the overall performance impact of $\mathcal{L}_{cal}$.
+We also conduct an ablation study on the impact of active LoRA experts in Appendix A.4.
 5.4.1 Incremental Weighting on Calibration Term
@@
 Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
-Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintains similar accuracy.
-Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
-ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
-UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
-ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
-UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
+Table 4: Performance of UQ4CT with varying $\beta$ values on the OBQA dataset.
+Prediction accuracy and uncertainty calibration improve with increasing $\beta$, highlighting the effectiveness of $\mathcal{L}_{\mathrm{cal}}$.
+$\beta$ ACC $\uparrow$ ECE $\downarrow$
+0 $75.5_{2.91}$ $12.9_{1.99}$
+0.2 $76.0_{0.6}$ $7.47_{0.78}$
+0.5 $75.9_{0.31}$ $7.82_{0.93}$
+0.8 $76.6_{0.4}$ $5.94_{1.16}$
+1 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
+5.4.2 Sensitivity Test on Calibration Term
+To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14.
+This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities.
+We evaluate $\beta$ values of $0,0.2,0.5,0.8$ and $1$, where $\beta=0$ resembles the original MixLoRA method.
+Results in Table 4 demonstrate the effectiveness of the calibration loss.
+When $\beta=0$, the model is optimized without calibration on parameter mixtures, resulting in high ECE value.
+Even with small $\beta=0.2$ or $\beta=0.5$, the ECE scores drastically improved compared to no calibration setting.
+Finally, when $\beta=1$, the calibration term effectively optimizes the conditional parameter mixtures to generate outputs that fit data distribution well, resulting in lower ECE scores and higher accuracies.
 6 Discussion & Conclusion
@@
 The limitation of UQ4CT lies in its dependency on predictive correctness.
-For general language modeling tasks such as chat completion, there lacks a clear metric on whether the response is correct or not.
+For general language modeling tasks such as chat completion, there lacks a clear metric on response correctness.
 This limits the application of UQ4CT as naively token matching is a poor indicator of semantic correctness due to the ambiguous nature of language.
-For future work, we plan to continue exploring ways to adapt UQ4CT on open-ended problems that lacks a definitive metric.
-Reproducibility Statement
-We share our experimental details in Appendix A.1, and also provide the code and model weights for running experiments in the supplementary materials to reproduce our model performance results.
+For future work, we are exploring ways to adapt UQ4CT to open-ended problems that lack a definitive optimization objective.
 Acknowledgement
@@
 Appendix A Appendix
-A.1 Training Details
+A.1 Load Balancing Loss
+We follow the load balancing loss in [Li et al., 2024]. Given $N$ experts indexed by $i=1$ to $N$ and a batch $B$ with $T$ tokens, the auxiliary loss is computed as:
+$\mathcal{L}_{aux}=a\cdot N\cdot\sum^{N}_{i=1}\mathcal{F}_{i}\cdot\mathcal{P}_{i},$ (16)
+where
+$\mathcal{F}_{i}=\frac{1}{T}\sum_{x\in B}\mathbbm{1}\{argmax_{k}\mathcal{R}(x)_{k}=i\},\mathcal{P}_{i}=\frac{1}{T}\sum_{x\in B}\mathcal{R}(x)_{i}.$ (17)
+Here, $\mathcal{R}(\cdot)$ is the top-k router, $\mathcal{F}_{i}$ is the fraction of tokens dispatched to expert $i$ and $\mathcal{P}_{i}$ is the fraction of the router probability allocated for expert $i$.
+The final loss is multiplied by the expert count $N$ to keep the loss constant as the number of experts varies, and the constant term $a$ is set to $10^{-2}$as a multiplicative coefficient, which is large enough to ensure load balancing while remaining small enough not to overwhelm the primary objective.
+A.2 Experimental Results with Mistral-7B
+In this section, we present the results using Mistral-7B [Jiang et al., 2023], a different decoder-based LLM backbone.
+Table 5 shows the results of fine-tuning Mistral-7B on $4$ common-sense reasoning tasks and one domain-specific climate question-answering task.
+For each of the tasks, UQ4CT effectively calibrates the parameter mixtures, leading to the best ECE performance in $4$ out of $5$ tasks. This indicates the robustness of UQ4CT across different LLMs.
+Table 5: Performance comparison of different methods fine-tuned with Mistral-7B across $4$ common sense reasoning tasks and a domain-specific task.
+UQ4CT shows significant ECE improvements while maintaining high accuracy.
+Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
+ACC $\uparrow$ LoRA $70.3_{0.62}$ $84.8_{0.47}$ $70.2_{0.84}$ $82.8_{0.62}$ $72.5_{1.6}$
+MC Drop $69.6_{1.07}$ $84.6_{0.91}$ $69.6_{0.76}$ $82.6_{0.71}$ $72.5_{1.6}$
+Ensemble $71.8_{1.29}$ $84.2_{0.66}$ $71.0_{1.41}$ $82.5_{0.6}$ $72.9_{2.88}$
+LA $70.7_{1.82}$ $82.4_{2.05}$ $68.5_{3.31}$ $82.5_{0.77}$ $71.6_{1.56}$
+MixLoRA $73.1_{0.38}$ $85.5_{1.27}$ $71.2_{1.75}$ $83.3_{1.14}$ $72.0_{1.69}$
+UQ4CT $73.6_{0.28}$ $85.9_{0.82}$ $74.4_{0.82}$ $83.7_{1.22}$ $73.2_{1.29}$
+ECE $\downarrow$ LoRA $10.17_{0.24}$ $9.46_{1.62}$ $18.42_{1.91}$ $13.3_{0.25}$ $13.72_{2.62}$
+MC Drop $10.62_{0.51}$ $8.91_{1.35}$ $18.38_{1.66}$ $13.3_{0.31}$ $13.72_{2.61}$
+Ensemble $8.72_{1.13}$ $8.72_{1.49}$ $17.0_{0.97}$ $9.14_{2.82}$ $12.86_{1.78}$
+LA $5.33_{2.16}$ $20.3_{5.7}$ $21.27_{4.15}$ $\mathbf{6.41_{3.22}}$ $14.64_{2.21}$
+MixLoRA $8.81_{1.03}$ $8.16_{0.99}$ $15.51_{3.86}$ $10.53_{1.73}$ $14.05_{3.09}$
+UQ4CT $\mathbf{3.07_{0.83}}$ $\mathbf{5.7_{0.69}}$ $\mathbf{7.04_{0.58}}$ $\underline{7.92_{1.14}}$ $\mathbf{11.4_{1.14}}$
+A.3 Incremental Weighting on Calibration Term
+Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data.
+Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration.
+We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
+$\beta=\min\left\{1,\frac{\mathrm{current\_grad\_step}}{50}\right\}.$ (18)
+We choose $50$ gradient steps from our observation that training loss generally stabilizes after $50$ gradient steps, indicating the LoRA experts have learned some functional relationships from data.
+As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks.
+This demonstrates the advantage of uncertainty calibration even in the early stage.
+In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty.
+Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
+A.4 Deciding Number of Active Experts
+One important aspect of the LoRA MoE architecture is how many experts to activate.
+Here, we investigate the performance impact of different number of active LoRA experts.
+We evaluate the model performance with $1$ to $5$ active experts with $8$ in total.
+Table 6: Performance comparison of UQ4CT with varying number of experts on OBQA dataset. Top-2 expert selection strategy grants best accuracy and calibration.
+Top-K ACC $\uparrow$ ECE $\downarrow$
+Top-1 $74.8_{0.62}$ $7.69_{1.96}$
+Top-2 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
+Top-3 $75.2_{0.8}$ $5.8_{0.81}$
+Top-4 $75.8_{0.53}$ $7.67_{0.46}$
+Top-5 $75.3_{0.5}$ $6.29_{0.61}$
+As shown in Table 6, $2$ active experts give the optimal performance in terms of accuracy and ECE scores.
+One expert alone cannot capture complicated functional relationships, while more than 2 experts could potentially introduce redundant functional bases to the model, which deviates the output distribution more from data distribution, thus worsening predictive and calibration performance.
+Additionally, more active experts lead to a more flattened distribution across experts, which hardens the alignment of parameter mixtures during fine-tuning.
+A.5 Training Details
 We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability.
@@
 For Laplace-LoRA, we follow the Laplace hyperparameters in this Github Repository.
-A.2 Expected Calibration Error
+A.6 Expected Calibration Error
 Expected calibration error (ECE) is a commonly used metric to asses uncertainty quantification performance.
@@
 $\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}P(\hat{y}_{i}).$ (21)
-A.3 MMLU Distribution Shift Dataset Composition
+A.7 MMLU Distribution Shift Dataset Composition
   • Computer Science (CS):

### v3

@@
   1.
-Accurate uncertainty quantification of large language models (LLMs) provides credibility measure over their outputs.
-However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions due to the limitations in the models’ ability to generalize with limited data.
-Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage and fall short of calibrating epistemic uncertainty.
-To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates epistemic uncertainty over the space of functions that map input prompts to outputs.
+Accurate uncertainty quantification in large language models (LLMs) is essential for providing credible confidence estimates over their outputs.
+However, fine-tuned LLMs often exhibit overconfidence in uncertain predictions, which stems from their limited ability to generalize with sparse data.
+Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage, and thus fail to address the core issue: limited specialization of PEFT adapters to accurately capture task-specific input-output relationships.
+To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates uncertainty over the space of functions that map input prompts to outputs.
 We implement UQ4CT during the fine-tuning stage via a mixture-of-experts framework that hierarchically decomposes the functional space.
-We demonstrate that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks.
+Empirically, UQ4CT achieves over $25\%$ reduction in Expected Calibration Error (ECE) while preserving high accuracy across five benchmarks.
 Even under distribution shift, UQ4CT maintains superior ECE performance with high accuracy, showcasing improved generalizability.
 11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint.
-Under Review.
 Correspondence to: Yian Ma <yianma@ucsd.edu>.
 1 Introduction
-Quantifying credibility of the outputs has been one of the most important problems around large language models (LLMs) (Chang et al., 2024).
-In particular, fine-tuned LLMs often struggle with overconfidence in gauging the uncertain of their outputs—oftentimes due to limited training data—and does not reflect the true credibility of the answers that they generate (Xiao et al., 2022a; Zhou et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023).
-Such overconfidence has become a crucial problem in safety-related decision making and in fields where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022).
-Methods that enhance uncertainty quantification of fine-tuned LLMs are sought after for the assurance of trustworthy predictions in applications.
-One salient challenge of uncertainty quantification in large models is the trade-off among accuracy, calibration, and tractability.
-One would first like to maintain the capability of the LLMs and not sacrifice accuracy of the generated outputs.
-Given this desideratum, a number of existing works focus on perturbing the prompts that are fed into the language models and quantify the uncertainty associated with such perturbations (Hou et al., 2023; Farquhar et al., 2024; Gao et al., 2024).
-These methods capture the aleatoric uncertainty assuming that the trained model corresponds to the ground-truth data distribution.
-However, these approaches cannot capture the epistemic uncertainty that arises from the model being trained on limited data.
-To capture this epistemic uncertainty, one would need to introduce changes or perturbations to the model in accordance with the data distribution.
-On the tractability side, it is not realistic to consider perturbations to all the model parameters for an LLM (Minaee et al., 2024).
-We therefore propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the epistemic uncertainty in LLMs.
-Since we only focus on the epistemic uncertainty over the functional space that maps input prompts to the outputs, we no longer need to consider uncertainty in the entire parameter space.
-Instead, we only need to measure the uncertainty regarding the functional relationships.
-Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the interesting functional relationships arising from the underlying data.
-In particular, we make use of the low rank adaptation (LoRA, Hu et al.
-(2021a)) ensemble at each layer of the LLM to form the basis functions.
-We then utilize the mixture of experts (MoE) architecture (Li et al., 2024) to hierarchically combine the basis functions and form the functional space.
-During fine-tuning, UQ4CT simultaneously learns the parameters in the LoRA experts and the prompt-dependent mixture weights to output the calibrated distribution over the functional space.
-During inference, the LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input.
-The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space.
-We calibrate functional level epistemic uncertainty to align with predictive correctness during training time.
-This approach significantly improves uncertainty estimations of the model on its predictions without compromising accuracy.
+Quantifying the credibility of outputs has been one of the most important problems around large language models (LLMs)(Chang et al., 2024).
+In particular, fine-tuned LLMs often struggle with overconfidence in their outputs due to limited training data, failing to reflect the true credibility of their answers(Xiao et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023).
+Such overconfidence can assert misinformation with high certainty, making it difficult for users to discern truth from falsehood.
+This has become a crucial problem in safety-critical decision making and scientific domains where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022).
+Methods that enhance uncertainty quantification for fine-tuned LLMs are therefore essential to ensure trustworthy predictions.
+One salient challenge of uncertainty quantification in large language models is the trade-off among accuracy, calibration, and efficiency.
+Ideally, one seeks to calibrate model uncertainty without degrading accuracy or slowing output generation.
+Recent approaches often focus on prompt perturbation—modifying the model input and quantifying the resulting prediction variance (Hou et al., 2023; Gao et al., 2024)—or sampling multiple completions to measure prediction disagreement (Farquhar et al., 2024).
+However, these methods generally assume the model is already well-aligned with the data distribution, and thus struggle to capture uncertainty arising during fine-tuning, especially the generalization gap due to adaptation on limited data.
+Additionally, since these methods require multiple forward passes per input, they incur significant computational overhead, limiting scalability.
+Beyond prompt-level approaches, Bayesian methods and ensemble-based uncertainty quantification have been established for fine-tuned LLMs, often in conjunction with low-rank adaptation (LoRA)  (Hu et al., 2021a).
+Methods such as Monte Carlo dropout (Gal and Ghahramani, 2016), checkpoint ensembles (Chen et al., 2017), deep ensembles (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023), and Laplace-LoRA (Yang et al., 2024a) apply Bayesian inference or ensembling over LoRA parameters to capture uncertainty arising from model adaptation and limited data.
+While Bayesian and ensemble methods estimate uncertainty after fine-tuning by analyzing the learned parameter space, they do not address the limitations caused by sparse data during fine-tuning.
+This post hoc perspective can miss uncertainty that arises when adapting to new tasks with limited data.
+Figure 1: Left: The Mixture of Experts (MoE) architecture captures diverse functional relationships by dynamically routing inputs to different expert modules based on the prompt.
+Right: The standard LoRA approach lacks such functional-space diversity, limiting its capacity to capture variations in the functional relationships during fine-tuning.
+To overcome this, we shift focus from parameter-space to functional-space uncertainty quantification.
+The functional space encompasses the input-output mappings the model can realize, capturing the true variability in its predictions.
+By calibrating uncertainty at this level during fine-tuning, we ensure the model’s confidence better reflects its actual predictive reliability.
+We therefore introduce Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), a method that explicitly calibrates functional-level uncertainty in LLMs during fine-tuning.
+UQ4CT leverages ensembles of LoRA modules at each layer to construct a rich set of basis functions.
+We then employ a Mixture-of-Experts (MoE) architecture (Li et al., 2024) to hierarchically combine these basis functions, forming a flexible functional space (see Figure 1).
+During fine-tuning, UQ4CT jointly learns the LoRA expert parameters and calibrates the prompt-dependent function mixture to align functional-level uncertainty with predictive correctness, enabling the model to output calibrated distributions over the functional space.
+During inference, LoRA experts offer diverse functional relationships acquired during fine-tuning, while MoE routers dynamically select the most relevant experts for each input prompt.
+This selection is guided by functional-level uncertainty calibration performed throughout fine-tuning, which aims to optimize the choice of the correct functional relationship for each prompt.
+More accurate expert selection enables the model to learn diverse functional relationships.
+As a result, the model’s uncertainty estimates become better aligned, enhancing calibration without compromising accuracy.
 To summarize, our contributions include:
-  • A novel epistemic uncertainty quantification approach with MoE architecture during fine-tuning to quantify functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
-  • A novel calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
-  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance under distribution shifts on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
+  • A novel uncertainty quantification approach for LLMs with MoE architecture during fine-tuning to quantify functional-level uncertainty and align with the probability of predictive correctness, which mitigates overconfidence issue and improves generalizability.
+  • A new calibration loss that incorporates predictive correctness probability to dynamically align the prompt-dependent LoRA mixture for better uncertainty estimation.
+  • Hierarchical decomposition of functional-level uncertainty into layer-wise mixture weights with guarantee that our calibration loss aligns mixture weights with predictive correctness.
+  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance without compromising accuracy under distribution shift on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks.
 2 Preliminaries
-2.1 Low-rank Adaptation (LoRA)
+Low-rank Adaptation (LoRA).
 LLMs have numerous large weight matrices to perform matrix multiplication, denoted as $\mathbf{W}_{0}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}}$ that maps inputs $\mathbf{x}$ to outputs $\mathbf{h}$.
@@
 LoRA greatly mitigates this memory cost as the tripled memory consumption only applies to LoRA adapters.
-2.2 Mixture of Experts (MoE)
+Mixture of Experts (MoE).
 LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is an efficient approach to scale the number of parameters while maintaining the same computational bounds.
@@
 The router is a linear layer that maps the input hidden state $\mathbf{h}$ to a probability distribution of candidate experts.
-The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and feed forward layer (FFN), together with residual connections.
-Formally, given $\mathbf{h}_{o}=[s_{0},s_{1},...,s_{n}]$ as the tokenized input text, the output of $\ell$-th layer is generated as:
+The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and the feedforward layer (FFN), together with residual connections.
+Formally, given $\mathbf{h}^{1}$ as the tokenized input text, the output of $\ell$-th layer is generated as:
 $\mathbf{z}^{\ell}=f_{attn}(\text{FFN}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{FFN}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (2)
 Here, $f_{attn}$ represents the attention calculation step.
-Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension.
-With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-n gate router $R(\cdot)$ chooses n experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
-$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-n}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (3)
+Let $\mathbf{h}^{\ell}\in\mathbb{R}^{1\times d}$ $(1\leq\ell\leq L)$ denote the output hidden state at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension.
+With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router $\tilde{R}(\cdot)$ chooses $k$ experts with highest probability given a hidden state $\mathbf{h}^{\ell}$:
+$\tilde{R}^{\ell}(\mathbf{h}^{\ell})=\text{Keep-Top-k}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}^{\ell})).$ (3)
 Finally, we obtain the final MixLoRA prediction with:
-$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
-where $\mathbf{W}$ is the pretrained weights of the target fine-tuning layer and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
+$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}\tilde{R}^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}_{pre}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
+where $\mathbf{W}^{\ell}_{pre}$ is the frozen pretrained weight at layer $\ell$ and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
 With MixLoRA defined in Equation 4, we can apply MixLoRA layers at $q,k,v$ encoding and FFN layers:
 $\mathbf{z}^{\ell}=f_{attn}(\text{MixLoRA}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{MixLoRA}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (5)
-Note that the calculation of $E^{\ell}_{k}$ involves $\mathbf{W}^{\ell}$, which represents the weight of the original ATTN or FFN layer.
-2.3 Aleatoric and Epistemic Uncertainties
-In machine learning models, uncertainty can be categorized into aleatoric (data-wise) and epistemic (model-wise) uncertainty (Hora, 1996; Hüllermeier and Waegeman, 2021).
-For LLMs, aleatoric uncertainty arises from the inherently ambiguous and context dependent nature of natural languages where a single phrase or sentence can have multiple valid interpretations in different contexts.
-Epistemic uncertainty is introduced by the model’s lack of knowledge due to limited learning capabilities, suboptimal modeling or sparse training data.
-Epistemic uncertainty is highly related to several well-known limitations of generative models.
-For example, it has been observed that when an LLM is pretrained on a diverse range of text data, it is generally well-calibrated, i.e.
-the predicted probability of the next token generally aligns with what is observed in real text.
-However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a).
-A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target task (Lin et al., 2023; Luo et al., 2023).
-Motivated by these observations, we explore functional-level epistemic uncertainty in generative models and aim to develop metrics that assess model performance on specific problem instances to fine-tune the parameter mixture.
 3 Methodology
-The high level goal of UQ4CT is to balance the exploration and exploitation of different LoRA experts during fine-tuning.
-In particular, we incorporate the functional-level epistemic uncertainty (FEU) to calibrate the prompt-dependent parameter mixture with LoRA MoE.
-Assume that our answer $a$ is generated via a mixture of mechanisms or models $M$, conditioning on the input prompt $x$.
-Assume that $e(a)$ is an embedding of $a$ so that least squares distance is a natural distance on the space of $e(a)$.
-For an expressive enough model class $\mathcal{M}$, and a calibrated distribution $P(M|x)$ over the model class, we can measure the deviation of the generated answer to the “ideal” one $a_{*}=f_{*}(x)$ as:
-$\mathbb{E}_{M\sim P(M|x)}\left[\mathbb{E}_{a\sim P(a|M,x)}\left[\|e(a)-e(a_{*})\|^{2}\right]\right]$
-$=\mathbb{E}_{P(M|x)}\underbrace{\left[\mathbb{E}_{P(a|M,x)}\left[\|e(a)-\mathbb{E}_{P(a|M,x)}[e(a)]\|^{2}\right]\right]}_{\text{Aleatoric Uncertainty}}+\underbrace{\mathbb{E}_{P(M|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-e(a_{*})\|^{2}\right]}_{\text{Epistemic Uncertainty}}.$ (6)
-Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$.
-We hence name the task “functional-level uncertainty quantification”.
-3.1 Functional-level Epistemic Uncertainty
-Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty.
-Experts $B_{1...K}A_{1...K}$ capture functional relationships in the data throughout fine-tuning, the weights $\alpha_{1...K}$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$.
-In the UQ4CT workflow, we align this uncertainty with predictive correctness.
-When the router makes a correct prediction, the loss reinforces this decision, thereby increasing the router confidence in its selection, which aligns with a lower epistemic uncertainty.
-Conversely, when the router makes an incorrect prediction, the loss penalizes this selection, potentially causing the router to distribute its probabilities more broadly across experts, which is indicative of higher epistemic uncertainty.
-Motivated by the decomposition in Eq.
-(3) for least squares loss, we consider a general distance $d$, and formally define epistemic uncertainty as the variance caused by model training procedure.
-Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs.
-Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
-$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(\cdot|x)}\mathbb{E}_{M^{\prime}\sim P(\cdot|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right]$ (7)
-Here, $a^{\prime}$ and $a$ represents the outputs sampled from the parameter mixtures $M^{\prime}$ and $M$ conditioned on the prompt $x$.
-The epistemic uncertainty measures the least squares distance between $e(a)$ and $e(a^{\prime})$.
+The high-level goal of UQ4CT is to leverage the ensemble of prompt-dependent LoRA mixture-of-experts (MoE) to guide and calibrate the confidence of the model during fine-tuning.
+By quantifying the variability in how different LoRA experts are combined for each input, UQ4CT enables the model to adaptively select expert mixtures that reflect the true uncertainty in its predictions.
+Our approach not only encourages the model to exploit confident expert combinations for accurate predictions but also promotes exploration of alternative experts when uncertainty is high, ultimately leading to better-calibrated and more reliable model outputs.
+Figure 2: MoE architecture to capture functional-level uncertainty.
+LoRA experts ($B_{i}$, $A_{i}$) capture diverse functional bases, while the top-$k$ router assigns mixture weights based on the input hidden state.
+The calibration loss aligns functional-level uncertainty with prediction correctness, encouraging confident expert selection for correct predictions and higher uncertainty for incorrect ones.
 3.1 Decomposition of the Functional Space
-Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the functional relationships in the underlying data.
-Then the functional space can be decomposed into a mixture of the LoRA ensembles.
-Naively, this decomposition writes:
-$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{1}=1}^{K}\sum_{k^{2}=1}^{K}\cdots\sum_{k^{L}=1}^{K}\alpha_{k^{1},k^{2},\dots,k^{L}}g_{k^{L}}^{L}\left(g_{k^{L-1}}^{L-1}\left(\dots g_{k^{1}}^{1}(\mathbf{x})\right)\right),$ (8)
-$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right),$ (9)
-where $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e.
-layer norm, attention calculation, etc.).
-This linear decomposition requires examining $K^{L}$ parameters to span the entire functional space.
-We instead consider a hierarchical decomposition of $f(\mathbf{x})$:
-$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\Bigg{(}\sum_{k^{L-1}=1}^{K}\alpha^{L-1}_{k^{L-1}}g_{k^{L-1}}^{L-1}\Bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$
-Here the output function is determined by the combinations of the weights $\alpha^{\ell}_{k^{\ell}}$ across all the layers.
-As a result, we only need $K\cdot L$ parameters to represent various compositions of the functional relationships.
-To further improve the efficiency of representation, we apply a sparse dynamic routing and choose the mixture weights $\alpha_{k^{\ell}}=R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}$ that depend on the input from the previous layer:
-$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
-3.3 Quantifying FEU with LoRA MoE Framework
-We quantify the functional-level epistemic uncertainty (FEU) with the MoE architecture, which is represented by the embedding $e(a)$ in Eq.
-(7).
-As shown in Figure 1, the LoRA experts that are made of matrices $B^{1...L}_{1...K},A^{1...L}_{1...K}$ in Eq.
-(1) capture important functional relationships in the data during fine-tuning.
-We treat these functional relationships represented by the LoRA experts as basis functions.
-As shown in Eq.
-(10), conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions, hierarchically through each layer.
-Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions.
-As shown in Figure 1, the weights $\alpha_{k}^{\ell}$ from the top-n router dynamically mix these basis functions given the input hidden state.
-The top-n weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection.
-With mixture of LoRA experts architecture, we follow routing mechanisms of the MoE layers as in Eq.
-(3) and (4).
-Specifically, we employ top-2 gate routers, which chooses $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
-$\vspace{-0.5mm}R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (11)
-Given an input prompt $x$ with length $s$, we model functional-level epistemic uncertainty (FEU) by aggregating $R^{\ell}(\mathbf{h}^{\ell}_{i})$ over both layer dimension and sequence dimension:
-$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right].$ (12)
-3.4 Training Calibration Loss
-The FEU model of the epistemic uncertainty enables us to calibrate the parameter mixtures against predictive accuracy for better aligned output distributions.
-Specifically for the MoE top-k routers, we design the following calibration loss for training:
-$\mathcal{L}_{\mathrm{cal}}=\left\|\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FEU}(x)\right\|^{2},$ (13)
-where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$.
-We employ a one-hot definition of the ground truth confidence.
-When the prediction from current mixture model matches the ground truth, the ground truth confidence is $1$.
-Otherwise, when the predictions do not match, the ground truth confidence is 00.
-3.5 Calibrating the Mixture
-Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration with functional-level epistemic uncertainty.
-As shown in Figure 1, $\mathcal{L}_{\mathrm{cal}}$ effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness.
-Ideally, when the $K$ LoRA experts together capture all the functional relationships across the data distribution with cross entropy during fine-tuning, our proposed loss $\mathcal{L}_{cal}$ also finds proper mixture of LoRA experts conditioned on the input $x$ by conditionally promoting expert exploitation and exploration.
-This allows the model to select correct functional relationships regarding $x$ to generate an output that better matches the data distribution, which grants calibrated uncertainty estimations.
-Load balancing is a common regularization to ensure even utilization across experts with the MoE architecture (Fedus et al., 2022).
-We follow the load balancing loss $\mathcal{L}_{b}$ proposed by Li et al.
-(2024) and define our loss function as:
-$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
-where $\mathbf{CE}$ represents cross entropy loss, $\gamma$ and $\beta$ are the hyperparameters of two auxiliary terms. Details about $\mathcal{L}_{b}$ can be found in Appendix A.1.
+Given the immense size of both the pre-training dataset and the model, we posit that the pretrained network contains submodules capable of expressing a wide range of functional relationships present in the data.
+During fine-tuning, our focus is on the functional space spanned by the model, which can be effectively captured as a mixture of LoRA experts, each representing a distinct basis function.
+However, naively decomposing the full functional space by considering all possible combinations of these experts quickly leads to a combinatorial explosion.
+Naive Decomposition.
+Denote the input prompt as $\mathbf{x}$ and the functional map from input to output as $f$.
+A straightforward approach is to express the function as a sum over all possible compositions of $K$ submodules per layer, across $L$ layers:
+$f(\mathbf{x})=\sum_{k^{1},\cdots,k^{L}=1}^{K}\alpha_{k^{1},\dots,k^{L}}g_{k^{L}}^{L}\bigg{(}\cdots\bigg{(}g_{k^{l}}^{l}\big{(}\dots g_{k^{1}}^{1}(\mathbf{x})\big{)}\bigg{)}\bigg{)},$ (6)
+where each $g_{k^{\ell}}^{\ell}$ denotes a particular variant (e.g., LoRA-adapted) of the $\ell$-th block and
+$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=\mathbf{h}^{\ell}+E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)},$ (7)
+here $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e.
+layer norm, attention calculation, etc.) and as defined in Eq.
+4, $E_{k^{\ell}}^{\ell}$ denotes the parameterized adaptation associated with the $k^{\ell}$-th expert in the $\ell$-th layer.
+However, this naive decomposition is intractable in practice, as it requires keeping track of $K^{L}$ mixture weights $\alpha_{k^{1},\ldots,k^{L}}$.
+This is an exponential growth in the number of parameters with respect to both the number of layers $L$ and the number of submodules per layer $K$, which makes the direct approach computationally infeasible for realistic network sizes.
+Hierarchical Decomposition.
+To address the combinatorial explosion of mixture weights, we instead propose a hierarchical decomposition.
+Here, the mixture at each layer is formed independently, and the output of each layer is a weighted sum over its submodules, with the weights themselves being layer-specific:
+$f(\mathbf{x})=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\bigg{(}\cdots\Bigg{(}\sum_{k^{l}=1}^{K}\alpha^{l}_{k^{l}}g_{k^{l}}^{l}\bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\bigg{)}\Bigg{)}\Bigg{)}.$ (8)
+Instead of needing $K^{L}$ mixture weights, this hierarchical structure only requires $K\cdot L$ weights $\alpha^{\ell}_{k^{\ell}}$, one for each submodule in each layer.
+This dramatically reduces the parameters required and makes the decomposition tractable, while still enabling a rich set of compositional functions.
+Dynamic, Input-Dependent Routing.
+To further enhance expressivity and efficiency, we allow the mixture weights to depend dynamically on the input at each layer.
+Specifically, we set the mixture weights to be a sparse routing function $R^{\ell}$ of the hidden state $\mathbf{h}^{\ell}$ at each layer, where $\alpha_{k^{\ell}}^{\ell}=\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=R_{k^{\ell}}^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell})).$
+Substituting this definition into the hierarchical mixture, the overall function $f(\mathbf{x})$ becomes:
+$f(\mathbf{x})=\sum_{k^{L}=1}^{K}R_{k^{L}}^{L}(f_{trans}^{L}(\mathbf{h}^{L}))g_{k^{L}}^{L}\Bigg{(}\dots\Bigg{(}\sum_{k^{1}=1}^{K}R_{k^{1}}^{1}(f_{trans}^{1}(\mathbf{h}^{1}))g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$ (9)
+This recursive formulation shows that at each layer, the submodules are weighted according to the input-dependent routing function, which adapts based on the current hidden state.
+For a layer-wise perspective, the computation at each layer $\ell$ can be explicitly written as:
+$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R_{k^{\ell}}^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
+Here, $R^{\ell}$ selects the most relevant submodules for a given input, allowing the network to adaptively compose its computation path at each layer in a sparse fashion.
+3.2 Quantifying FLU with MixLoRA
+In the previous section, we have established a parsimonious representation of the functional space.
+In this section, we choose a simple function to encode the functional level uncertainty and provide the intuition as follows.
+We first note that given a fixed MoE model architecture, the larger weight a mixture component has, the more certain we are about that component.
+The model uncertainty of the MixLoRA architecture is quantified by considering perturbations $\Delta f(\mathbf{x})$ to the model $f(\mathbf{x})$.
+Following the discussion and notation in Sec.
+3.1, we can show that these perturbations are instantiated in the space of the mixture weights $\alpha$ (as defined in Eq.
+8 and 9):
+Fact 3.1 (Model Perturbation Structure).
+Under regularity assumptions on the residual connection architecture, perturbations $\Delta f(x)$ to the model $f(x)$ approximately decompose as:
+$\Delta f(x)\approx\sum_{\ell=1}^{L}\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell}).$
+In high-dimensional settings, the basis functions are approximately orthogonal.
+Thus, the perturbation $\Delta f(x)$ is entirely represented by the set $\left\{\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\right\}_{k^{\ell}=1,\dots,K}^{\ell=1,\dots,L}$.
+Therefore, the functional-level uncertainty (FLU) can be generally modeled as a linear function over the mixture weights:
+$\text{FLU}(x)=U\left(\left\{\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\right\}_{k^{\ell}=1,\dots,K}^{\ell=1,\dots,L}\right),$
+where $U(\cdot)$ denotes a linear aggregation function. Details of derivation is presented in Appendix A.1.
+In practice, as illustrated in Figure 2, the top-$k$ router at each layer produces a sparse probability vector $\alpha^{\ell}=(\alpha_{1}^{\ell},...,\alpha_{k}^{\ell})$, dynamically mixing the basis functions captured by the LoRA experts given the current hidden state.
+The values of the top-$K$ routing weights that contribute to the final output hidden state serve as a direct quantification of the model’s uncertainty at functional-level.
+We follow the routing mechanisms used in MoE layers (see Eq.
+(3) and (4)), employing top-$2$ gate routers for the mixture.
+At each layer $\ell$, we compute the raw router probabilities and retrain the largest two:
+$\tilde{R}^{\ell}(\mathbf{h}^{\ell})=\text{Keep-Top-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}^{\ell})).$ (11)
+Given an input prompt $x$ of length $s$, we aggregate the router weights over the selected experts and across all layers to estimate the FLU:
+$\text{FLU}(x)=\frac{1}{L}\sum_{\ell=1}^{L}\sum_{i=1}^{2}\tilde{R}_{i}^{\ell}(\mathbf{h}^{\ell}).$ (12)
+This formulation provides an efficient approach to quantify functional-level uncertainty in LoRA MoE architectures.
+3.3 Calibration Loss
+The FLU model provides a principled way to calibrate the mixture parameters against predictive accuracy, enabling better alignment between output distributions and true model confidence.
+Specifically, for MoE top-$k$ routers, we design the following calibration loss for training:
+$\mathcal{L}_{\mathrm{cal}}=\left(\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FLU}(x)\right)^{2}.$ (13)
+The first term is an indicator function that equals $1$ if the model prediction matches the ground truth $y^{*}$ for prompt $x$, and 00 otherwise, corresponding to a one-hot definition of ground truth confidence.
+This calibration loss directly encourages the functional-level uncertainty (FLU) to reflect the true correctness of the model’s predictions.
+As shown in Figure 2, when the mixture model predicts correctly, the loss pushes FLU toward $1$ (high confidence); when incorrect, toward 00 (low confidence).
+In other words, when optimized over the data distribution, the calibration loss pushes FLU to represent the probability of prediction correctness.
+We formally state this property as follows and present the proof in Appendix A.1:
+Proposition 3.2 (Truthfulness of Calibration Loss).
+Let the calibration risk be defined as the expectation of the calibration loss over the data distribution:
+¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
+If this calibration risk is optimized over the data distribution $\mathcal{D}$, then the optimal solution is $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$; that is, the optimally trained FLU corresponds to the probability that the model’s prediction is correct.
+3.4 Fine-tuning with the Total Loss
+Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration according to functional-level uncertainty.
+As shown in Figure 2, $\mathcal{L}_{\mathrm{cal}}$ aligns uncertainty with predictive correctness, increasing the router probability of the selected expert for correct predictions (exploitation) and decreasing it for incorrect ones (exploration).
+Ideally, when the $K$ LoRA experts collectively capture the relevant functional relationships in the data during fine-tuning via cross-entropy loss, $\mathcal{L}_{\mathrm{cal}}$ further guides the model to select appropriate mixtures of LoRA experts conditioned on the input $\mathbf{x}$.
+This targeted selection enables the model to match the data distribution more closely and provides more calibrated uncertainty estimates.
+To ensure balanced expert utilization, we incorporate a load balancing loss $\mathcal{L}_{b}$ as proposed by Li et al. (2024). Our overall loss function is:
+$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}b+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
+where $\mathbf{CE}$ is the cross-entropy loss, and $\gamma$, $\beta$ are hyperparameters for the auxiliary terms. Details of $\mathcal{L}_{b}$ are provided in Appendix A.2.
 4 Related Work
-4.1 Parameter-Efficient Fine Tuning for LLMs
-Large Language Models (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks.
-Building on these advances, instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems (Wu et al., 2023b; Achiam et al., 2023b).
-However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
-To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT) (Mangrulkar et al., 2022), model distillation (Liu et al., 2023; Xiao et al., 2023), quantization (Frantar et al., 2022; Xiao et al., 2022b), and pruning (Ma et al., 2023).
-Among these, LoRA (Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference.
-For example, VeRA (Kopiczko et al., 2023) introduces learnable scaling vectors to modify shared pairs of frozen random matrices across layers, while FedPara (Hyeon-Woo et al., 2021) focuses on low-rank Hadamard products for federated learning settings.
-Tied-Lora (Renduchintala et al., 2023) applies weight tying to further minimize the number of trainable parameters.
-AdaLoRA (Zhang et al., 2023) uses Singular Value Decomposition (SVD) to prune less important singular values for efficient updates, and DoRA (Liu et al., 2024) separates pre-trained weights into magnitude and direction components, applying LoRA to update the directional component during fine-tuning, thus reducing the number of parameters to be trained.
-4.2 Mixture of Experts
-The Mixture-of-Experts concept (Jacobs et al., 1991; Wang et al., 2020), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data.
-Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
-Different MoE architectures have since emerged, distinguished by their expert sampling and routing strategies.
-For example, LLaVA-MoLE (Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines.
-Other MoE-based approaches include MoRAL (Yang et al., 2024a), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE (Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting.
-PESC (Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational costs.
-MoE-LoRA (Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE (Gou et al., 2023) activates task-specific model parameters based on instruction clusters.
-MixLoRA (Li et al., 2024) implements a high-throughput framework for LoRA MoE training and inference process, constructing LoRAs as stochastic experts to reduce computational overhead while expanding model capacity.
-Despite the performance improvements these architecture advancements have brought, the overconfidence problem of fine-tuned models is lacking attention (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023).
-Enhancing the uncertainty estimation capabilities of these models is fundamental toward more reliable, interpretable and trustworthy applications of LLMs.
-4.3 Uncertainty Quantification in LLMs
-Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks (Gal and Ghahramani, 2015, 2016a; Malinin and Gales, 2018; Ovadia et al., 2019; Dusenberry et al., 2020; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023).
-This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research (Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023).
-LLMs, particularly in generative tasks, pose unique challenges, especially when it comes to measuring the uncertainty of their outputs (Liu et al., 2019; Malinin and Gales, 2021; Kuhn et al., 2023; Lin et al., 2023).
-The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs (Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
-Established uncertainty quantification methods have been studied in conjunction with the LoRA structure.
-Monte-Carlo dropout (Gal and Ghahramani, 2016b) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them.
+Mixture of LoRA Experts.
+Large Language Models (LLMs) have achieved impressive performance across diverse NLP tasks (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c), with instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) further boosting their adaptability for conversational AI (Wu et al., 2023b; Achiam et al., 2023).
+However, scaling LLMs increases the resource demands of full fine-tuning.
+Parameter-efficient fine-tuning (PEFT) methods (Mangrulkar et al., 2022)—such as LoRA (Hu et al., 2021b) and its variants (Kopiczko et al., 2023; Hyeon-Woo et al., 2021; Renduchintala et al., 2023; Zhang et al., 2023; Liu et al., 2024)—reduce adaptation costs by updating a subset of parameters.
+Recent advances combine PEFT with the Mixture-of-Experts (MoE) framework (Jacobs et al., 1991; Wang et al., 2020), which sparsely activates expert subnetworks for greater model capacity and specialization.
+MoE-based LLMs leverage expert routing and parameter-efficient adaptations to target new domains or tasks efficiently.
+Notably, methods such as MoRAL (Yang et al., 2024b), LoRAMoE (Dou et al., 2024), PESC (Wu et al., 2024b), MoE-LoRA (Luo et al., 2024), and MixLoRA (Li et al., 2024) optimize domain-specific routing, mitigate forgetting, and enable scalable, high-throughput training and inference with mixtures of LoRA experts.
+Uncertainty Quantification in LLMs.
+Established uncertainty quantification methods have been studied in conjunction with the LoRA structure for LLMs.
+Monte-Carlo dropout (Gal and Ghahramani, 2016) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them.
 Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty.
 Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration.
-Laplace-LoRA (Yang et al., 2024b) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
-However, none of these methods provide calibrations regarding epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
+Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
+Prompt-perturbation and resampling-based approaches have also been explored for uncertainty quantification in LLMs.
+These methods estimate uncertainty by measuring prediction variability under different prompt formulations or sampled input variants, without altering model parameters (Farquhar et al., 2024; Hou et al., 2023; Gao et al., 2024).
+This line of work leverages the inherent sensitivity of LLMs to input perturbations as a means to assess model confidence, providing a complementary perspective to parameter-based methods.
+Ye et al.
+(2024) benchmark LLMs using conformal prediction, which quantifies uncertainty by constructing prediction sets with guaranteed coverage, where set size directly reflects model uncertainty.
 5 Experiments
-5.1 Datasets
-We include $5$ multiple-choice question answering benchmarks: OpenBookQA (OBQA, Mihaylov et al.
-(2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science.
-We also use computer science, law, medication and engineering subsets from MMLU dataset (Hendrycks et al., 2020) to evaluate performance under distribution shift.
-We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance.
-We report the average model performance over $3$ random runs and the standard deviations in the subscript.
-Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task.
+Datasets.
+We evaluate on five multiple-choice QA benchmarks: OpenBookQA (OBQA) (Mihaylov et al., 2018), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) (Clark et al., 2018), BOOLQ (Clark et al., 2019), and ClimateQA—a domain-specific climate science benchmark.
+To assess robustness under distribution shift, we ensemble the domain-specific MMLU subtasks (Hendrycks et al., 2020) into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health.
+Details for the ensemble are provided in Appendix A.8.
+Models are fine-tuned on the public training split and evaluated on the test split for each benchmark.
+Figure 3: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task.
 UQ4CT shows substantial ECE improvements while maintaining high accuracy.
-Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
-ACC $\uparrow$ LoRA $69.5_{1.93}$ $74.8_{1.39}$ $53.8_{0.6}$ $72.1_{0.87}$ $59.9_{2.13}$
-MC Drop $66.8_{3.66}$ $76.8_{1.30}$ $50.9_{2.01}$ $74.8_{1.34}$ $58.2_{2.11}$
-Ensemble $66.2_{3.7}$ $71.2_{1.0}$ $47.5_{0.57}$ $75.5_{1.4}$ $59.6_{6.9}$
-LA $68.7_{1.32}$ $74.6_{2.11}$ $51.4_{0.83}$ $70.8_{1.24}$ $55.2_{3.29}$
-MixLoRA $71.5_{1.05}$ $77.7_{2.27}$ $54.3_{1.07}$ $75.5_{2.91}$ $61.6_{1.76}$
-UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
-ECE $\downarrow$ LoRA $11.9_{0.78}$ $11.9_{2.04}$ $19.4_{4.75}$ $10.2_{1.07}$ $14.3_{0.64}$
-MC Drop $12.2_{0.85}$ $11.9_{1.99}$ $19.8_{4.85}$ $10.9_{0.24}$ $14.3_{0.56}$
-Ensemble $7.28_{2.3}$ $9.1_{1.49}$ $10.23_{1.39}$ $8.83_{2.35}$ $13.5_{3.29}$
-LA $17.1_{1.72}$ $16.6_{3.7}$ $18.1_{0.5}$ $17.2_{1.2}$ $12.6_{1.9}$
-MixLoRA $7.88_{2.09}$ $9.09_{0.81}$ $10.74_{1.07}$ $12.9_{1.99}$ $12.5_{1.32}$
-UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
-5.2 Experiment Setup
-We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024) and compare the average performance in three random runs and report the mean and standard deviation with following baselines.
-We use the publicly available LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model.
-In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-fowrward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer).
-Details are provided in Appendix A.5
+Experiment Setup.
+We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024).
+We use the LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model.
+In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-forward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer).
+Details are provided in Appendix A.6.
   • LoRA (Hu et al., 2021a). We use standard LoRA fine-tuning as lower performance bound.
-  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016b) keeps dropout on at both training and testing time.
+  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016) keeps dropout on at both training and testing time.
 By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions.
@@
 We combine deep ensemble with LoRA by fine-tuning $3$ randomly initialized LoRAs together and ensembling their output as final predictions.
-  • Laplace-LoRA (LA) (Yang et al., 2024b) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
+  • Laplace-LoRA (LA) (Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
   • MixLoRA (Li et al., 2024) incorporates LoRAs via a router network to reduce computational overhead while expanding model capacity.
 We add this as a baseline to resemble plain LoRA MoE model performance.
-Evaluation Metrics.
-We measure the prediction accuracy on the validation set for all $5$ tasks.
-For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al.
-(2017)) with $15$ bins, which measures the alignment between predicted probabilities and empirical accuracy.
-We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications.
-Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts.
-Metric details are given in Appendix A.6.
-5.3 Results
-We assess the prediction accuracy and uncertainty calibration of models under both in-distribution and distribution shift scenarios.
-The in-distribution scenario examines the alignment of the fine-tuned model on the target downstream task, while the distribution shift scenario evaluates the generalizability of the model on novel tasks beyond the fine-tuned domain.
-These two scenarios combined provides a comprehensive assessment of model robustness in real-world applications where it is essential for the model to excel on its primary task while maintaining the ability to effectively handle unforeseen or out-of-distribution inputs.
+Evaluation.
+We evaluate prediction accuracy on the validation set across all five tasks.
+For uncertainty calibration, we use the expected calibration error (ECE; Guo et al.
+(2017); more details in A.7) to measure the alignment between predicted probabilities and actual outcomes.
+To assess robustness under distribution shifts, we fine-tune models on the OBQA dataset and evaluate them following (Yang et al., 2024a).
+We use ARC-C and ARC-E to represent smaller distribution shifts, as these datasets focus on general science reasoning similar to OBQA but are more challenging and diverse.
+For larger shifts, we utilize the four aforementioned domain-specific MMLU subtasks, which span a wide range of expertise from elementary to professional levels.
+This domain specificity represents a greater distribution shift from the general common sense focus of OBQA.
+The in-distribution scenario tests model alignment on the target task, while the distribution shift scenario assesses generalizability to novel domains.
+Together, they provide a comprehensive evaluation for real-world applications, ensuring strong performance on the primary task and resilience to out-of-distribution inputs.
 Table 1: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA2-7B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks.
@@
 MC Drop $74.8_{1.34}$ $58.7_{2.07}$ $66.6_{3.30}$ $36.0_{1.69}$ $30.3_{2.25}$ $35.1_{0.86}$ $39.1_{1.35}$
-Ensemble $75.5_{1.4}$ $57.7_{0.78}$ $69.1_{0.48}$ $36.7_{2.18}$ $30.3_{1.13}$ $35.3_{1.02}$ $39.9_{1.89}$
+Ensemble $75.5_{1.4}$ $57.7_{0.78}$ $69.1_{0.48}$ $\mathbf{36.7_{2.18}}$ $30.3_{1.13}$ $35.3_{1.02}$ $39.9_{1.89}$
 LA $70.8_{1.24}$ $58.7_{0.58}$ $67.9_{0.41}$ $33.7_{1.22}$ $29.6_{1.32}$ $35.4_{0.75}$ $38.5_{1.61}$
-MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $69.2_{1.02}$ $35.2_{2.92}$ $30.3_{0.98}$ $35.9_{0.43}$ $40.6_{1.13}$
-UQ4CT $77.3_{1.36}$ $58.8_{1.06}$ $65.8_{1.31}$ $36.2_{1.24}$ $34.1_{2.31}$ $35.8_{1.01}$ $40.0_{1.24}$
+MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $\mathbf{69.2_{1.02}}$ $35.2_{2.92}$ $30.3_{0.98}$ $\mathbf{35.9_{0.43}}$ $\mathbf{40.6_{1.13}}$
+UQ4CT $\mathbf{77.3_{1.36}}$ $\mathbf{58.8_{1.06}}$ $65.8_{1.31}$ $\underline{36.2_{1.24}}$ $\mathbf{34.1_{2.31}}$ $\underline{35.8_{1.01}}$ $\underline{40.0_{1.24}}$
 ECE $\downarrow$ LoRA $10.2_{1.07}$ $16.7_{2.28}$ $13.3_{2.48}$ $29.7_{2.69}$ $32.3_{1.85}$ $29.2_{3.08}$ $31.0_{2.13}$
@@
 UQ4CT $\mathbf{5.0_{1.15}}$ $\mathbf{8.9_{3.46}}$ $\mathbf{6.5_{1.85}}$ $\mathbf{19.6_{2.90}}$ $\mathbf{23.1_{1.17}}$ $\mathbf{25.9_{3.43}}$ $\mathbf{21.9_{3.49}}$
-5.3.1 In-distribution Performance
-We first evaluate UQ4CT and baseline models fine-tuned on the $4$ common sense reasoning tasks and the climate question answering task under the in-distribution scenario, where models are trained and evaluated on different splits of the same dataset.
-Note that one of the key advantages of UQ4CT is that uncertainty calibration happens during the fine-tuning stage with little computational overhead, while other UQ methods require repetitive sampling or other post-hoc complexities.
-As shown in Table 1, UQ4CT demonstrates notable improvements in uncertainty calibration across a variety of tasks.
-Across all benchmarks, UQ4CT maintains competitive accuracy (ACC) compared to the baseline methods.
-For example, on the BoolQ and ClimateQA datasets, UQ4CT achieves accuracy rates of $73.5\%$ and $63.3\%$, respectively.
-This empirically demonstrates that UQ4CT is capable of maintaining high accuracy with uncertainty calibration, which assures the gain in UQ performance does not compromise accuracy.
-The most substantial performance improvement is observed in the reduction of Expected Calibration Error (ECE).
-UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks.
-Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration.
-In addition to experiments on LLaMA-2-7B in the main text, we also present additional experiments on fine-tuning Mistral-7B in Appendix A.2 for more comprehensive evaluation of our method.
-For both LLaMA-2-7B and Mistral-7B models, UQ4CT consistently shows substantial improvements in uncertainty calibration.
-The improvements are critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
-5.3.2 Performance Under Distribution Shift
+5.1 In-distribution Performance
+As shown in Figure 3, UQ4CT achieves notable gains in uncertainty calibration across diverse tasks, while maintaining competitive accuracy (ACC) relative to baseline approaches.
+For instance, on BoolQ and ClimateQA tasks, UQ4CT attains accuracy rates of $73.5\%$ and $63.3\%$, demonstrating that improved uncertainty quantification does not come at the expense of predictive performance.
+The most significant improvements are seen in reduced Expected Calibration Error (ECE).
+UQ4CT consistently lowers ECE by over $25\%$ on average across benchmarks, and unlike other approaches, it continues to perform well even on challenging datasets such as ARC-C, achieving an ECE of $6.1$.
+To further validate our method, we include results from fine-tuning both LLaMA-2-7B (main text) and Mistral-7B (Appendix A.3).
+Across both models, UQ4CT delivers substantial and consistent improvements in uncertainty calibration.
+These results underscore the practical value of UQ4CT, particularly in scenarios where reliable uncertainty estimates are crucial, such as safety-critical applications.
+A key advantage of UQ4CT is that it incorporates uncertainty calibration directly during fine-tuning, incurring minimal computational overhead compared to other uncertainty quantification (UQ) methods, which often require costly repetitive sampling or post-hoc adjustments.
+5.2 Performance Under Distribution Shift
 Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a).
 Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios.
-Similar to the dataset setup in (Yang et al., 2024b), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA.
-For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health.
-These tasks have very broad coverage of the domain task at various knowledge levels ranging from elementary school to professionals.
-This domain-specificity demonstrates larger distribution shift from OBQA, which is a general common sense reasoning task.
-Details of the ensemble is provided in Appendix.
-A.7.
-The distribution shift evaluations are provided in Table 2.
-UQ4CT provides substantial improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios.
-For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario.
-For the more challenging larger distribution shifts, UQ4CT still achieves the best ECE performance among all baseline models.
-Note that UQ4CT also achieves competitive prediction accuracy across all domain-specific tasks.
-This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
-Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintaining similar accuracy.
-Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
-ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
-UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
-ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
-UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
-5.4 Ablation Study
-In this section, we conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$.
-We first evaluate the incremental weighting performance of the calibration term, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning.
-Then, we perform a sensitivity test, where we explore the overall performance impact of $\mathcal{L}_{cal}$.
+Table 1 presents the distribution shift evaluations.
+UQ4CT achieves substantial ECE improvements while maintaining high accuracy across both smaller and larger distribution shifts.
+For smaller shifts, UQ4CT’s ECE remains comparable to the in-distribution scenario.
+Under larger shifts, UQ4CT attains the lowest ECE among all baselines and delivers competitive accuracy on all domain-specific tasks.
+These results demonstrate that aligning uncertainty at the functional level with predictive correctness improves generalizability and mitigates overconfidence in fine-tuned models.
+5.3 Ablation Studies
+We conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$.
+We first perform a sensitivity test, in which we explore the impact of $\mathcal{L}_{cal}$ on the overall performance.
+Then we evaluate the incremental weighting performance of the calibration term in Appendix A.4, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning.
 We also conduct an ablation study on the impact of active LoRA experts in Appendix A.5.
-5.4.1 Incremental Weighting on Calibration Term
-Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data.
-Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration.
-We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
-$\beta=\min\left\{1,\frac{\mathrm{current\_grad\_step}}{50}\right\}.$ (15)
-We choose $50$ gradient steps from our observation that training loss generally stabilizes after $50$ gradient steps, indicating the LoRA experts have learned some functional relationships from data.
-As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks.
-This demonstrates the advantage of uncertainty calibration even in the early stage.
-In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty.
-Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
 Table 2: Performance of UQ4CT with varying $\beta$ values on the OBQA dataset.
 Prediction accuracy and uncertainty calibration improve with increasing $\beta$, highlighting the effectiveness of $\mathcal{L}_{\mathrm{cal}}$.
-$\beta$ ACC $\uparrow$ ECE $\downarrow$
-0 $75.5_{2.91}$ $12.9_{1.99}$
-0.2 $76.0_{0.6}$ $7.47_{0.78}$
-0.5 $75.9_{0.31}$ $7.82_{0.93}$
-0.8 $76.6_{0.4}$ $5.94_{1.16}$
-1 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
-5.4.2 Sensitivity Test on Calibration Term
+$\beta$ 00 $0.2$ $0.5$ $0.8$ $1$ $1.2$ $1.5$ $1.8$ $2$
+ACC $\uparrow$ $75.5_{2.91}$ $76.0_{0.6}$ $75.9_{0.31}$ $76.6_{0.4}$ $77.3_{1.36}$ $\mathbf{77.9_{0.6}}$ $77.2_{0.91}$ $76.7_{0.82}$ $76.6_{0.92}$
+ECE $\downarrow$ $12.9_{1.99}$ $7.47_{0.78}$ $7.82_{0.93}$ $5.94_{1.16}$ $\mathbf{5.0_{1.15}}$ $6.49_{2.65}$ $9.19_{0.86}$ $8.11_{1.43}$ $6.65_{1.85}$
+Sensitivity Test on Calibration Term.
 To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14.
 This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities.
-We evaluate $\beta$ values of $0,0.2,0.5,0.8$ and $1$, where $\beta=0$ resembles the original MixLoRA method.
+We evaluate $\beta$ values ranging from 00 to $2$, where $\beta=0$ resembles the original MixLoRA method.
 Results in Table 2 demonstrate the effectiveness of the calibration loss.
@@
 In this work, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which addresses the overconfidence issues commonly encountered during fine-tuning of large language models.
-We present a functional perspective on quantifying epistemic uncertainty in LLMs and utilize it for uncertainty-calibrated fine-tuning.
-By incorporating functional-level epistemic uncertainty quantification with a mixture-of-experts framework, our proposed uncertainty-calibrated training loss effectively addresses the challenge of overconfidence in fine-tuned LLMs by significantly improving uncertainty calibration while maintaining high accuracy.
+We present a functional perspective on quantifying uncertainty in LLMs and utilize it for uncertainty-calibrated fine-tuning.
+By incorporating functional-level uncertainty quantification with a mixture-of-experts framework, our proposed uncertainty-calibrated training loss effectively addresses the challenge of overconfidence in fine-tuned LLMs by significantly improving uncertainty calibration while maintaining high accuracy.
 Our evaluations demonstrate that UQ4CT reduces the Expected Calibration Error by more than 25% without compromising accuracy across a variety of downstream tasks, including common-sense and domain-specific reasoning, under in-distribution and out-of-distribution scenarios.
@@
 Appendix A Appendix
-A.1 Load Balancing Loss
+A.1 Theoretical Derivation of the Method
+In this section, we provide complete theoretical statements and proofs that are used in Sec. 3 to derive our method.
+Fact A.1 (Model Perturbation Structure, Restatement of Fact 3.1).
+Assume that in the residual connection architecture in each layer: $g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=\mathbf{h}^{\ell}+E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)}$, the Lipschitzness of the residual term $E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)}$ is much smaller as compared to that of $\mathbf{h}^{\ell}$ itself: $\left\|E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\hat{\mathbf{h}}^{\ell})\big{)}-E_{k^{\ell}}^{\ell}\big{(}f_{trans}^{\ell}(\mathbf{h}^{\ell})\big{)}\right\|=o\left(\|\hat{\mathbf{h}}^{\ell}-\mathbf{h}^{\ell}\|\right)$.
+Under this regularity assumption, perturbations $\Delta f(x)$ to the model $f(x)$ approximately decomposes as:
+$\Delta f(x)\approx\sum_{\ell=1}^{L}\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell}).$
+Proof of Fact 3.1 and A.1.
+In each layer, we can decompose the perturbation to the output as follows:
+$\Delta\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}\left[\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})+\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot\Delta g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\right],\quad\forall\ell=1,\dots,L,$
+where the input $\mathbf{h}^{1}=x$ and the output $f(x)=\mathbf{h}^{L+1}$.
+Due to the residual connection architecture and our assumption on the regularity of the residual term, we have:
+$\Delta g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=\Delta\mathbf{h}^{\ell}+\Delta E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right)=\Delta\mathbf{h}^{\ell}+o\left(\Delta\mathbf{h}^{\ell}\right).$
+Hence,
+$\Delta\mathbf{h}^{\ell+1}\approx\sum_{k^{\ell}=1}^{K}\left(\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})+\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot\Delta\mathbf{h}^{\ell}\right)=\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})+\Delta\mathbf{h}^{\ell}.$
+Expanding this recursion, the output perturbation can be approximated as:
+$\Delta f(x)=\Delta\mathbf{h}^{L+1}\approx\sum_{\ell=1}^{L}\sum_{k^{\ell}=1}^{K}\Delta\alpha_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})\cdot g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell}).$
+Proposition A.2 (Calibration Loss, Restatement of Proposition 3.2).
+Let the calibration risk be defined as the expectation of the calibration loss over the data distribution:
+¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
+If this calibration risk is optimized over the data distribution $\mathcal{D}$, then the optimal solution is $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$; that is, the optimally trained FLU corresponds to the probability that the model’s prediction is correct.
+Proof of Proposition 3.2 and A.2.
+Expanding the calibration risk, we have:
+¯$\bar{\mathcal{L}}_{\mathrm{cal}}$ $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}$
+$=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left[\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}\right]^{2}$
+$\quad-2\,\mathbb{E}_{x}\mathbb{E}_{y^{*}|x}\left[\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}\cdot\mathrm{FLU}(x)\right]+\mathbb{E}_{x}\left[\mathrm{FLU}(x)^{2}\right]$
+$=C+\mathbb{E}_{x}\left[\mathrm{FLU}(x)^{2}-2\,\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))\cdot\mathrm{FLU}(x)\right],$
+where $C$ is a constant independent of $\mathrm{FLU}(x)$. Thus, the calibration risk is minimized when $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$. ∎
+A.2 Load Balancing Loss
 We follow the load balancing loss in [Li et al., 2024]. Given $N$ experts indexed by $i=1$ to $N$ and a batch $B$ with $T$ tokens, the auxiliary loss is computed as:
@@
 The final loss is multiplied by the expert count $N$ to keep the loss constant as the number of experts varies, and the constant term $a$ is set to $10^{-2}$as a multiplicative coefficient, which is large enough to ensure load balancing while remaining small enough not to overwhelm the primary objective.
-A.2 Experimental Results with Mistral-7B
+A.3 Experimental Results with Mistral-7B
 In this section, we present the results using Mistral-7B [Jiang et al., 2023], a different decoder-based LLM backbone.
@@
 For each of the tasks, UQ4CT effectively calibrates the parameter mixtures, leading to the best ECE performance in $4$ out of $5$ tasks. This indicates the robustness of UQ4CT across different LLMs.
+Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintains similar accuracy.
+Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
+ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
+UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
+ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
+UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
 Table 4: Performance comparison of different methods fine-tuned with Mistral-7B across $4$ common sense reasoning tasks and a domain-specific task.
@@
 MixLoRA $73.1_{0.38}$ $85.5_{1.27}$ $71.2_{1.75}$ $83.3_{1.14}$ $72.0_{1.69}$
-UQ4CT $73.6_{0.28}$ $85.9_{0.82}$ $74.4_{0.82}$ $83.7_{1.22}$ $73.2_{1.29}$
+UQ4CT $\mathbf{73.6_{0.28}}$ $\mathbf{85.9_{0.82}}$ $\mathbf{74.4_{0.82}}$ $\mathbf{83.7_{1.22}}$ $\mathbf{73.2_{1.29}}$
 ECE $\downarrow$ LoRA $10.17_{0.24}$ $9.46_{1.62}$ $18.42_{1.91}$ $13.3_{0.25}$ $13.72_{2.62}$
@@
 UQ4CT $\mathbf{3.07_{0.83}}$ $\mathbf{5.7_{0.69}}$ $\mathbf{7.04_{0.58}}$ $\underline{7.92_{1.14}}$ $\mathbf{11.4_{1.14}}$
-A.3 Incremental Weighting on Calibration Term
+A.4 Incremental Weighting on Calibration Term
 Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data.
@@
 This demonstrates the advantage of uncertainty calibration even in the early stage.
-In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty.
+In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high functional-level uncertainty.
 Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
-A.4 Deciding Number of Active Experts
+A.5 Deciding Number of Active Experts
 One important aspect of the LoRA MoE architecture is how many experts to activate.
@@
 Additionally, more active experts lead to a more flattened distribution across experts, which hardens the alignment of parameter mixtures during fine-tuning.
-A.5 Training Details
+A.6 Training Details
 We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability.
@@
 We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training.
+Our model is trained on A100 GPU, with 20GB GPU memory consumption per task.
+Training time is from $25$ to $50$ minutes depending on the task.
 The experimental setup for single LoRA based models is similar with LoRA ranks set to $80$ to accommodate the MoE model size.
@@
 For Laplace-LoRA, we follow the Laplace hyperparameters in this Github Repository.
-A.6 Expected Calibration Error
+A.7 Expected Calibration Error
 Expected calibration error (ECE) is a commonly used metric to asses uncertainty quantification performance.
@@
 $\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}P(\hat{y}_{i}).$ (20)
-A.7 MMLU Distribution Shift Dataset Composition
+In this paper, we use an ECE bin size of $15$, following the experiment setup in Laplace-LoRA [Yang et al., 2024a].
+A.8 MMLU Distribution Shift Dataset Composition
   • Computer Science (CS):
@@
     • Virology
+A.9 Prompt Perturbation Comparison
+Here, we compare our method with SPUQ [Gao et al., 2024], which perturbs the prompt, aggregates predictions and confidences to measure uncertainty.
+We test SPUQ and UQ4CT with LLama2-7b as the base model.
+As shown in Table 6, SPUQ’s large ECE values suggest that simply aggregating predictions from perturbed prompts does not adequately calibrate model confidence, highlighting the limitations of prompt perturbation as an uncertainty quantification strategy for LLMs, especially for smaller models.
+Table 6: Performance comparison of UQ4CT and SPUQ across five tasks. UQ4CT achieves higher accuracy and substantially lower ECE than SPUQ.
+Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
+ACC $\uparrow$ SPUQ $62.20$ $75.14$ $45.99$ $59.00$ $60.64$
+UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
+ECE $\downarrow$ SPUQ $30.47$ $13.13$ $24.74$ $20.00$ $19.98$
+UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$

### v4

@@
 Here, $\Delta\mathbf{W}$ is calculated as the product of two matrices, $\mathbf{B}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{lr}}}$ and $\mathbf{A}\in\mathbb{R}^{n_{\text{lr}}\times n_{\text{in}}}$ where $n_{\text{lr}}$ is significantly smaller than $n_{\text{in}}$ or $n_{\text{out}}$.
-For example, we use $n_{\text{lr}}=32$ while $n_{\text{in}}=n_{\text{out}}=4096$ for the Llama2-7b model (Touvron et al., 2023a).
+For example, we use $n_{\text{lr}}=32$ while $n_{\text{in}}=n_{\text{out}}=4096$ for the Llama3.1-8B model (Dubey et al., 2024).
 Therefore, the total number of LoRA parameters for this $\Delta\mathbf{W}$ is $n_{\text{lr}}(n_{\text{in}}+n_{\text{out}})$, which is far smaller than the parameter count of the full matrix, $n_{\text{in}}n_{\text{out}}$.
@@
 Here, $f_{attn}$ represents the attention calculation step.
-Let $\mathbf{h}^{\ell}\in\mathbb{R}^{1\times d}$ $(1\leq\ell\leq L)$ denote the output hidden state at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension.
-With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router $\tilde{R}(\cdot)$ chooses $k$ experts with highest probability given a hidden state $\mathbf{h}^{\ell}$:
+[…]
 $\tilde{R}^{\ell}(\mathbf{h}^{\ell})=\text{Keep-Top-k}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}^{\ell})).$ (3)
@@
 (3) and (4)), employing top-$2$ gate routers for the mixture.
-At each layer $\ell$, we compute the raw router probabilities and retrain the largest two:
+At each layer $\ell$, we compute the raw router probabilities and retain the largest two:
 $\tilde{R}^{\ell}(\mathbf{h}^{\ell})=\text{Keep-Top-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}^{\ell})).$ (11)
@@
 $\mathcal{L}_{\mathrm{cal}}=\left(\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FLU}(x)\right)^{2}.$ (13)
-The first term is an indicator function that equals $1$ if the model prediction matches the ground truth $y^{*}$ for prompt $x$, and 00 otherwise, corresponding to a one-hot definition of ground truth confidence.
+The first term is an indicator function that equals $1$ if the model prediction matches the ground truth $y^{*}$ for prompt $x$, and 0 otherwise, corresponding to a one-hot definition of ground truth confidence.
 This calibration loss directly encourages the functional-level uncertainty (FLU) to reflect the true correctness of the model’s predictions.
-As shown in Figure 2, when the mixture model predicts correctly, the loss pushes FLU toward $1$ (high confidence); when incorrect, toward 00 (low confidence).
+As shown in Figure 2, when the mixture model predicts correctly, the loss pushes FLU toward $1$ (high confidence); when incorrect, toward 0 (low confidence).
 In other words, when optimized over the data distribution, the calibration loss pushes FLU to represent the probability of prediction correctness.
@@
 Let the calibration risk be defined as the expectation of the calibration loss over the data distribution:
-¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
+$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
 If this calibration risk is optimized over the data distribution $\mathcal{D}$, then the optimal solution is $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$; that is, the optimally trained FLU corresponds to the probability that the model’s prediction is correct.
@@
 To ensure balanced expert utilization, we incorporate a load balancing loss $\mathcal{L}_{b}$ as proposed by Li et al. (2024). Our overall loss function is:
-$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}b+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
-where $\mathbf{CE}$ is the cross-entropy loss, and $\gamma$, $\beta$ are hyperparameters for the auxiliary terms. Details of $\mathcal{L}_{b}$ are provided in Appendix A.2.
+$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
+where $\mathbf{CE}$ is the cross-entropy loss, and $\gamma$, $\beta$ are hyperparameters for the auxiliary terms.
+We fix $\gamma$, $\beta$ to $1$ for our experiments.
+Details of $\mathcal{L}_{b}$ are provided in Appendix A.2.
 4 Related Work
 Mixture of LoRA Experts.
-Large Language Models (LLMs) have achieved impressive performance across diverse NLP tasks (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c), with instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) further boosting their adaptability for conversational AI (Wu et al., 2023b; Achiam et al., 2023).
+Large Language Models (LLMs) have achieved impressive performance across diverse NLP tasks (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023a, b), with instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) further boosting their adaptability for conversational AI (Wu et al., 2023b; Achiam et al., 2023).
 However, scaling LLMs increases the resource demands of full fine-tuning.
@@
 Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
+Bayesian Low-Rank Adaptation by Backpropagation (BLoB) (Wang et al., 2024) extends the LA method by jointly optimizing the mean and covariance of LoRA parameters via backpropagation throughout fine-tuning.
 Prompt-perturbation and resampling-based approaches have also been explored for uncertainty quantification in LLMs.
@@
 Models are fine-tuned on the public training split and evaluated on the test split for each benchmark.
-Figure 3: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task.
+Table 1: Performance comparison of different methods fine-tuned with LlaMA3.1-8B across four common sense reasoning tasks and a domain-specific task.
 UQ4CT shows substantial ECE improvements while maintaining high accuracy.
+Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
+ACC $\uparrow$ LoRA $\underline{89.73_{0.58}}$ $\underline{88.82_{1.82}}$ $78.21_{0.73}$ $88.00_{1.22}$ $78.25_{1.29}$
+MC Drop $9.65_{0.55} 88$$8.16_{1.75}$ $77.14_{0.69}$ $87.12_{1.18}$ $78.19_{1.35}$
+Ensemble $\mathbf{89.87_{0.43}}$ $\mathbf{89.14_{1.31}}$ $78.81_{0.96}$ $86.47_{0.42}$ $78.53_{2.98}$
+MixLoRA $88.68_{0.82}$ $87.74_{0.36}$ $78.56_{1.87}$ $\underline{88.27_{0.50}}$ $\underline{79.94_{1.29}}$
+LA $89.58_{0.19}$ $86.22_{3.52}$ $78.00_{3.76}$ $86.00_{6.01}$ $79.82_{3.48}$
+BLoB(Mean) $9.02_{0.93} 88$$8.71_{0.82}$ $79.37_{0.71}$ $87.60_{1.04}$ $79.02_{0.50}$
+BLoB(N=10) $89.39_{1.13}$ $87.96_{0.62}$ $\mathbf{80.08_{1.55}}$ $87.13_{0.88}$ $79.02_{0.50}$
+UQ4CT $9.17_{1.33} 88$$8.66_{0.20}$ $\underline{79.60_{1.21}}$ $\mathbf{88.40_{0.35}}$ $\mathbf{79.97_{0.85}}$
+ECE $\downarrow$ LoRA $15.82_{0.57}$ $6.55_{1.70}$ $14.07_{0.68}$ $7.30_{0.43}$ $13.70_{1.50}$
+MC Drop $14.73_{0.54}$ $6.48_{1.74}$ $14.12_{0.71}$ $7.24_{0.39}$ $13.11_{1.46}$
+Ensemble $14.56_{0.55}$ $7.08_{0.73}$ $13.71_{1.29}$ $8.63_{0.38}$ $14.69_{0.84}$
+MixLoRA $15.85_{0.76}$ $7.79_{0.45}$ $13.71_{1.90}$ $6.58_{0.21}$ $14.68_{0.09}$
+LA $3.78_{0.60}$ $7.63_{1.71}$ $8.92_{4.16}$ $1.97_{5.97} 11$$1.48_{1.66}$
+BLoB(Mean) $7.54_{0.57}$ $4.89_{0.32}$ $11.26_{1.13}$ $6.83_{0.90}$ $12.74_{0.88}$
+BLoB(N=10) $\underline{2.76_{0.41}}$ $\mathbf{3.35_{0.50}}$ $\underline{6.81_{1.43}}$ $\underline{3.84_{1.00}}$ $\underline{11.96_{2.57}}$
+UQ4CT $\mathbf{1.79_{0.43}}$ $\underline{3.97_{0.78}}$ $\mathbf{4.43_{0.82}}$ $\mathbf{3.34_{1.60}}$ $\mathbf{9.36_{2.77}}$
 Experiment Setup.
 We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024).
-We use the LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model.
-In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-forward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer).
+We use the Llama-3.1-8B (Touvron et al., 2023c) as our base model.
+In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-forward networks in LLaMA-3.1-8B (gate layer, down layer and up layer).
 Details are provided in Appendix A.6.
-  • LoRA (Hu et al., 2021a). We use standard LoRA fine-tuning as lower performance bound.
-  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016) keeps dropout on at both training and testing time.
-By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions.
-To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al.
-(2022).
-  • Deep Ensemble (Lakshminarayanan et al., 2017) averages the predictions from each ensemble member which have been trained with varying random initialization.
-We combine deep ensemble with LoRA by fine-tuning $3$ randomly initialized LoRAs together and ensembling their output as final predictions.
-  • Laplace-LoRA (LA) (Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
-  • MixLoRA (Li et al., 2024) incorporates LoRAs via a router network to reduce computational overhead while expanding model capacity.
-We add this as a baseline to resemble plain LoRA MoE model performance.
+Baselines.
+We compare UQ4CT with state-of-the-art uncertainty estimation methods along with naive fine-tuning applied to the LoRA adapters of LLMs, including LoRA (Hu et al., 2021a), Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016), Deep Ensemble (Lakshminarayanan et al., 2017), Laplace-LoRA (LA) (Yang et al., 2024a), Bayesian Low-Rank Adaptation by Backpropagation (BLoB) (Wang et al., 2024) and MixLoRA (Li et al., 2024).
+Note that for BLoB(N=10), the method performs $10$ forward passes with differently sampled LoRA parameters for each question, which is a unfair computational budget advantage compared against UQ4CT with only $1$ forward pass.
 Evaluation.
@@
 (2017); more details in A.7) to measure the alignment between predicted probabilities and actual outcomes.
-To assess robustness under distribution shifts, we fine-tune models on the OBQA dataset and evaluate them following (Yang et al., 2024a).
+To assess robustness under distribution shifts, we fine-tune models on the OBQA dataset and evaluate them following Yang et al.
+(2024a).
 We use ARC-C and ARC-E to represent smaller distribution shifts, as these datasets focus on general science reasoning similar to OBQA but are more challenging and diverse.
@@
 Together, they provide a comprehensive evaluation for real-world applications, ensuring strong performance on the primary task and resilience to out-of-distribution inputs.
-Table 1: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA2-7B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks.
+Table 2: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA3.1-8B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks.
 UQ4CT shows substantial ECE improvements while maintaining high accuracy.
@@
 Metrics Methods OBQA ARC-C ARC-E CS Eng Law Health
-ACC $\uparrow$ LoRA $72.1_{0.87}$ $58.6_{1.93}$ $66.5_{3.38}$ $35.5_{2.35}$ $30.8_{1.72}$ $34.9_{1.41}$ $39.1_{1.52}$
-MC Drop $74.8_{1.34}$ $58.7_{2.07}$ $66.6_{3.30}$ $36.0_{1.69}$ $30.3_{2.25}$ $35.1_{0.86}$ $39.1_{1.35}$
-Ensemble $75.5_{1.4}$ $57.7_{0.78}$ $69.1_{0.48}$ $\mathbf{36.7_{2.18}}$ $30.3_{1.13}$ $35.3_{1.02}$ $39.9_{1.89}$
-LA $70.8_{1.24}$ $58.7_{0.58}$ $67.9_{0.41}$ $33.7_{1.22}$ $29.6_{1.32}$ $35.4_{0.75}$ $38.5_{1.61}$
-MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $\mathbf{69.2_{1.02}}$ $35.2_{2.92}$ $30.3_{0.98}$ $\mathbf{35.9_{0.43}}$ $\mathbf{40.6_{1.13}}$
-UQ4CT $\mathbf{77.3_{1.36}}$ $\mathbf{58.8_{1.06}}$ $65.8_{1.31}$ $\underline{36.2_{1.24}}$ $\mathbf{34.1_{2.31}}$ $\underline{35.8_{1.01}}$ $\underline{40.0_{1.24}}$
-ECE $\downarrow$ LoRA $10.2_{1.07}$ $16.7_{2.28}$ $13.3_{2.48}$ $29.7_{2.69}$ $32.3_{1.85}$ $29.2_{3.08}$ $31.0_{2.13}$
-MC Drop $10.9_{0.24}$ $16.7_{2.20}$ $13.2_{2.21}$ $23.2_{2.32}$ $31.6_{1.64}$ $28.0_{2.93}$ $25.9_{2.27}$
-Ensemble $8.83_{2.35}$ $15.1_{1.09}$ $11.1_{0.99}$ $22.4_{1.32}$ $28.5_{2.13}$ $29.0_{1.37}$ $24.5_{0.39}$
-LA $17.2_{1.2}$ $16.2_{0.5}$ $24.4_{0.42}$ $28.6_{2.61}$ $30.5_{1.43}$ $29.5_{1.83}$ $30.7_{1.2}$
-MixLoRA $12.9_{1.99}$ $19.0_{1.88}$ $14.5_{2.57}$ $26.4_{3.25}$ $33.7_{1.87}$ $30.3_{2.27}$ $28.3_{1.07}$
-UQ4CT $\mathbf{5.0_{1.15}}$ $\mathbf{8.9_{3.46}}$ $\mathbf{6.5_{1.85}}$ $\mathbf{19.6_{2.90}}$ $\mathbf{23.1_{1.17}}$ $\mathbf{25.9_{3.43}}$ $\mathbf{21.9_{3.49}}$
+ACC $\uparrow$ LoRA $88.0_{1.22}$ $77.8_{0.16}$ $86.7_{0.77}$ $\mathbf{55.8_{0.52}}$ $54.3_{3.30}$ $44.9_{0.23}$ $58.8_{0.21}$
+MC Drop $87.1_{1.18}$ $77.1_{2.01}$ $86.9_{2.42}$ $54.4_{1.58}$ $54.1_{1.82}$ $45.0_{0.76}$ $58.3_{1.46}$
+Ensemble $86.5_{0.42}$ $78.2_{0.90}$ $85.4_{0.47}$ $53.8_{1.02}$ $52.4_{0.56}$ $45.0_{0.20}$ $60.6_{0.57}$
+MixLoRA $\underline{88.3_{0.50}}$ $78.1_{0.45}$ $86.7_{0.35}$ $53.1_{1.14}$ $\underline{54.7_{2.28}}$ $45.0_{1.46}$ $\underline{60.9_{1.04}}$
+LA $86.0_{6.01}$ $78.7_{0.55}$ $86.4_{0.76}$ $\underline{54.7_{1.82}}$ $53.6_{2.77}$ $44.9_{1.03}$ $59.7_{0.94}$
+BLoB(Mean) $87.6_{1.04}$ $\underline{79.5_{1.10}}$ $86.6_{0.65}$ $51.2_{0.99}$ $48.6_{1.44}$ $39.9_{7.85}$ $57.0_{3.51}$
+BLoB(N=10) $87.1_{0.88}$ $\mathbf{79.8_{1.06}}$ $\underline{87.2_{0.79}}$ $52.8_{1.28}$ $51.9_{3.13}$ $43.8_{4.60}$ $58.5_{5.33}$
+UQ4CT $\mathbf{88.4_{0.35}}$ $79.0_{0.56}$ $\mathbf{87.8_{0.47}}$ $53.3_{0.61}$ $\mathbf{61.1_{3.20}}$ $\mathbf{45.4_{0.50}}$ $\mathbf{61.1_{1.48}}$
+ECE $\downarrow$ LoRA $7.30_{0.43}$ $14.8_{0.62}$ $9.6_{0.69}$ $21.0_{3.05}$ $24.1_{3.63}$ $29.3_{1.98}$ $24.0_{1.81}$
+MC Drop $7.24_{0.39}$ $13.4_{2.15}$ $10.2_{1.89}$ $20.8_{3.26}$ $24.1_{0.77}$ $29.1_{0.64}$ $21.6_{3.92}$
+Ensemble $8.63_{0.38}$ $15.4_{0.46}$ $10.7_{0.56}$ $14.0_{3.18}$ $17.4_{1.98}$ $19.9_{2.95}$ $16.1_{2.07}$
+MixLoRA $6.58_{0.21}$ $14.5_{0.55}$ $9.9_{0.20}$ $17.1_{3.06}$ $17.8_{2.80}$ $21.6_{4.07}$ $18.0_{2.78}$
+LA $11.97_{5.97}$ $7.2_{0.5}$ $6.4_{0.42}$ $13.7_{2.14}$ $15.5_{2.0}$ $\underline{19.0_{0.54}}$ $15.7_{2.30}$
+BLoB(Mean) $6.83_{0.90}$ $11.37_{1.94}$ $6.6_{1.65}$ $17.2_{2.72}$ $18.5_{2.82}$ $22.6_{1.26}$ $16.9_{3.02}$
+BLoB(N=10) $\underline{3.84_{1.00}}$ $\underline{5.8_{0.96}}$ $\mathbf{3.0_{0.87}}$ $\underline{11.5_{2.76}}$ $\underline{14.9_{1.93}}$ $19.7_{3.21}$ $\underline{14.5_{3.38}}$
+UQ4CT $\mathbf{3.34_{1.60}}$ $\mathbf{3.6_{1.44}}$ $\underline{3.6_{1.32}}$ $\mathbf{10.8_{3.73}}$ $\mathbf{13.2_{1.86}}$ $\mathbf{18.1_{4.40}}$ $\mathbf{13.2_{4.06}}$
 5.1 In-distribution Performance
-As shown in Figure 3, UQ4CT achieves notable gains in uncertainty calibration across diverse tasks, while maintaining competitive accuracy (ACC) relative to baseline approaches.
-For instance, on BoolQ and ClimateQA tasks, UQ4CT attains accuracy rates of $73.5\%$ and $63.3\%$, demonstrating that improved uncertainty quantification does not come at the expense of predictive performance.
+As shown in Table 1, UQ4CT achieves notable gains in uncertainty calibration across diverse tasks, while maintaining competitive accuracy (ACC) relative to baseline approaches.
+For instance, on OBQA and ClimateQA tasks, UQ4CT attains accuracy rates of $88.4\%$ and $79.9\%$, demonstrating that improved uncertainty quantification does not come at the expense of predictive performance.
 The most significant improvements are seen in reduced Expected Calibration Error (ECE).
-UQ4CT consistently lowers ECE by over $25\%$ on average across benchmarks, and unlike other approaches, it continues to perform well even on challenging datasets such as ARC-C, achieving an ECE of $6.1$.
-To further validate our method, we include results from fine-tuning both LLaMA-2-7B (main text) and Mistral-7B (Appendix A.3).
+UQ4CT consistently lowers ECE by over $25\%$ on average across benchmarks, and unlike other approaches, it continues to perform well even on challenging datasets such as ARC-C, achieving an ECE of $4.4$.
+To further validate our method, we include results from fine-tuning both LLaMA-3.1-8B (main text) and Mistral-7B (Appendix A.3).
 Across both models, UQ4CT delivers substantial and consistent improvements in uncertainty calibration.
@@
 5.2 Performance Under Distribution Shift
-Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a).
+Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, c).
 Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios.
@@
 Prediction accuracy and uncertainty calibration improve with increasing $\beta$, highlighting the effectiveness of $\mathcal{L}_{\mathrm{cal}}$.
-$\beta$ 00 $0.2$ $0.5$ $0.8$ $1$ $1.2$ $1.5$ $1.8$ $2$
-ACC $\uparrow$ $75.5_{2.91}$ $76.0_{0.6}$ $75.9_{0.31}$ $76.6_{0.4}$ $77.3_{1.36}$ $\mathbf{77.9_{0.6}}$ $77.2_{0.91}$ $76.7_{0.82}$ $76.6_{0.92}$
-ECE $\downarrow$ $12.9_{1.99}$ $7.47_{0.78}$ $7.82_{0.93}$ $5.94_{1.16}$ $\mathbf{5.0_{1.15}}$ $6.49_{2.65}$ $9.19_{0.86}$ $8.11_{1.43}$ $6.65_{1.85}$
+$\beta$ 0 $0.2$ $0.5$ $0.8$ 11 $1.2$ $1.5$ $1.8$ 22
+[…]
+ECE $\downarrow$ $12.7_{1.92}$ $7.35_{0.75}$ $7.69_{0.89}$ $5.82_{1.12}$ $\mathbf{3.34_{1.60}}$ $6.31_{2.58}$ $9.03_{0.82}$ $7.96_{1.38}$ $6.52_{1.79}$
 Sensitivity Test on Calibration Term.
-To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14.
+To further understand the effectiveness of the calibration loss, we perform a sensitivity test of $\beta$ in Equation 14.
 This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities.
-We evaluate $\beta$ values ranging from 00 to $2$, where $\beta=0$ resembles the original MixLoRA method.
+We evaluate $\beta$ values ranging from 0 to $2$, where $\beta=0$ resembles the original MixLoRA method.
 Results in Table 3 demonstrate the effectiveness of the calibration loss.
@@
 Let the calibration risk be defined as the expectation of the calibration loss over the data distribution:
-¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
+$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
 If this calibration risk is optimized over the data distribution $\mathcal{D}$, then the optimal solution is $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$; that is, the optimally trained FLU corresponds to the probability that the model’s prediction is correct.
@@
 Expanding the calibration risk, we have:
-¯$\bar{\mathcal{L}}_{\mathrm{cal}}$ $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}$
+$\bar{\mathcal{L}}_{\mathrm{cal}}$ $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}$
 $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left[\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}\right]^{2}$
@@
 Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
-ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
-UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
-ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
-UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
-Table 4: Performance comparison of different methods fine-tuned with Mistral-7B across $4$ common sense reasoning tasks and a domain-specific task.
+ACC $\uparrow$ UQ4CT $9.17_{1.33} 88$$8.66_{0.20}$ $79.60_{1.21}$ $88.40_{0.35}$ $79.97_{0.85}$
+UQ4CT_Incremental $87.33_{0.24}$ $87.15_{0.95}$ $0.84_{1.03} 88$$8.53_{0.48}$ $75.87_{2.89}$
+ECE $\downarrow$ UQ4CT $\mathbf{1.79_{0.43}}$ $\mathbf{3.97_{0.78}}$ $\mathbf{4.43_{0.82}}$ $\mathbf{3.34_{1.60}}$ $\mathbf{9.36_{2.77}}$
+UQ4CT_Incremental $3.10_{0.18}$ $6.55_{1.42}$ $10.02_{1.95}$ $6.87_{1.68}$ $14.16_{0.91}$
+Table 5: Performance comparison of different methods fine-tuned with Mistral-7B across 44 common sense reasoning tasks and a domain-specific task.
 UQ4CT shows significant ECE improvements while maintaining high accuracy.
@@
 Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration.
-We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
+We conduct this study by incrementally increase $\beta$ from 0 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
 $\beta=\min\left\{1,\frac{\mathrm{current\_grad\_step}}{50}\right\}.$ (17)
@@
 Top-K ACC $\uparrow$ ECE $\downarrow$
-Top-1 $74.8_{0.62}$ $7.69_{1.96}$
-Top-2 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
-Top-3 $75.2_{0.8}$ $5.8_{0.81}$
-Top-4 $75.8_{0.53}$ $7.67_{0.46}$
-Top-5 $75.3_{0.5}$ $6.29_{0.61}$
+Top-1 $86.8_{0.59}$ $7.54_{1.89}$
+Top-2 $\mathbf{88.4_{0.35}}$ $\mathbf{3.34_{1.60}}$
+Top-3 $87.0_{0.77}$ $5.68_{0.78}$
+Top-4 $87.4_{0.51}$ $7.52_{0.44}$
+Top-5 $87.1_{0.48}$ $6.16_{0.58}$
 As shown in Table 6, $2$ active experts give the optimal performance in terms of accuracy and ECE scores.
@@
 We use batch size of $16$ to train our model.
-For climate task, we set the learning rate to $45e-$ 4 and dropout rate to $0.1$ to incorporate the small dataset size.
-For other tasks, we use $42e-$ 4 as our learning rate with dropout $0.05$.
+For climate task, we set the learning rate to 5e45e-4 and dropout rate to $0.1$ to incorporate the small dataset size.
+For other tasks, we use 2e42e-4 as our learning rate with dropout $0.05$.
 We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training.